### v1

# Precompactness in matrix weighted Bourgain-Morrey spaces

Abstract.
In this paper, we introduce matrix weighted Bourgain-Morrey spaces and obtain two sufficient conditions for precompact sets in matrix weighted Bourgain-Morrey spaces. We prove that the dyadic average operator is bounded on some matrix weighted Bourgain-Morrey spaces. With this result, we obtain the necessity for precompact sets in some matrix weighted Bourgain-Morrey spaces.
1 Introduction
In 1931, Kolmogorov [20] first discovered the characterization of precompact sets in $L^{p}([0,1])$ for $p\in(1,\infty)$. After that, there are many criteria for compactness of sets in Lebesgue spaces, which are called the Kolmogorov-Riesz compactness theorems on the Lebesgue spaces. More details and the history, we refer the reader to [14].
Inspired by [14], Clop-Cruz [7] obtained a compactness criterion in scalar weighted Lebesgue spaces $L^{p}(\omega)$ for $1<p<\infty$ with a scalar wight $\omega$ in Muckenhoupt class $A_{p}$. In [13], Guo and Zhao improved the result in [7] and obtained a compactness criterion in $L^{p}(\omega)$ for $p\in(0,\infty)$ with locally integrable weight $\omega$. In [21], Liu, Yang and Zhuo proved the Kolmogorov-Riesz compactness theorem in matrix weighted Lebesgue spaces $L^{p}(W)$ with $1<p<\infty$.
The theory of matrix weighted function spaces goes back to [32]. Indeed, in 1958, Wiener and Masani [32, Section 4] studied the matrix weighted $L^{2}(W)$ for the prediction theory for multivariate stochastic processes. In [29], Treil and Volberg introduced matrix class $\mathcal{A}_{2}$. Nazarov and Treil [23] and Volberg [30] extended $\mathcal{A}_{2}$ to $\mathcal{A}_{p}$ with $p\in(1,\infty)$. In [11], Goldberg showed that the matrix $\mathcal{A}_{p}$ condition leads to $L^{p}$ boundedness of a Hardy-Littlewood maximal function and obtained the boundedness of matrix weighted singular integral operators in Lebesgue spaces $L^{p},1<p<\infty$. In [24, 25, 26], Roudenko introduced the matrix-weighted homogeneous Besov spaces $\dot{B}^{s,q}_{p}(W)$ and matrix-weighted sequence Besov spaces $\dot{b}^{s,q}_{p}(W)$ and showed their equivalence via $\varphi$-transform and wavelets. In [9], Frazier and Roudenko introduced the matrix class $\mathcal{A}_{p},(0<p\leq 1)$ and studied the continuous and discrete matrix-weighted Besov spaces $\dot{B}^{s,q}_{p}(W)$ and $\dot{b}^{s,q}_{p}(W)$ with $0<p\leq 1$. In [10], Frazier and Roudenko introduced the homogeneous matrix-weighted Triebel-Lizorkin spaces $\dot{F}^{s,q}_{p}(W)$ for $s\in\mathbb{R},0<p<\infty,0<q\leq\infty$ and obtained the Littlewood-Paley characterizations of matrix-weighted Lebesgue spaces $L^{p}(W)$ and matrix-weighted Sobolev space $L^{p}_{k}(W)$ for $k\in\mathbb{N},1<p<\infty$. In [30], Volberg also introduced an analogue condition for the matrix class $\mathcal{A}_{\infty}$. In [3, 4, 5], Bu, Hytönen, Yang, Yuan studied the matrix weighted Besov-type and Triebel-Lizorkin-type spaces. Specifically, they introduced a new concept of the $\mathcal{A}_{p}$-dimension $\tilde{d}$, which is useful to the proof of main results of this paper. In [2], Bu, Hytönen, Yang, Yuan obtained several new characterizations of $\mathcal{A}_{p,\infty}$-matrix weights. For many other results on the matrix class $\mathcal{A}_{p}$ and matrix weighted function spaces, we refer the reader to [6, 8, 9, 10, 11, 24, 25, 31].
Bourgain [1] introduced a special case of Bourgain-Morrey spaces to study the Stein-Tomas (Strichartz) estimate. In [16], Hatano, Nogayama, Sawano, and Hakim researched the Bourgain-Morrey spaces from the viewpoints of harmonic analysis and functional analysis. In [19], Hu, Li and Yang introduced the Triebel-Lizorkin-Bourgain-Morrey spaces which connect Bourgain-Morrey spaces and global Morrey spaces.
Motivated by above literature, we will introduce matrix weighted Bourgain-Morrey spaces and research precompact sets in these spaces. The paper is organized as follows. In Section 2, dyadic cubes, the matrix class $\mathcal{A}_{p}$, $\mathcal{A}_{p}$-dimension $\tilde{d}$, matrix weighted Bourgain-Morrey spaces are given. The first result lies in Section 3. Specifically, a sufficient condition for totally bounded set in matrix weighted spaces $M_{p}^{t,r}(W)$ with $1\leq p<r<\infty$ or $1\leq p\leq t<r=\infty$ is obtained in Theorem 3.2. As a application, we get a criterion for totally bounded set in degenerate Bourgain-Morrey spaces with matrix weight. The second result (Theorem 4.1) is replacing the translation operator by the average operator. Note that the translation operator is not bounded on $L^{p}(W)$ and $M_{p}^{t,r}(W)$ in general. We prove that dyadic average operator is bounded on matrix weighted Bourgain-Morrey spaces with some conditions in Theorem 4.5. Using this result, we obtain the Kolmogorov-Riesz compactness theorem in matrix weighted Bourgain-Morrey spaces.
Throughout this paper, we let $c,C$ denote constants that are independent of the main parameters involved but whose value may differ from line to line. Let =$\mathbb{N}=\{1,2,3,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\ldots\}$. Let $\mathbb{Z}$ be the set of integers. Let $\chi_{E}$ be the characteristic function of the set $E\subset{\mathbb{R}^{n}}$. By $A\lesssim B$ we mean that $A\leq CB$ with some positive constant $C$ independent of appropriate quantities. By $A\approx B$, we mean that $A\lesssim B$ and $B\lesssim A$.
2 Preliminaries
For $j\in\mathbb{Z}$, $m\in\mathbb{Z}^{n}$, let $Q_{j,m}:=\prod_{i=1}^{n}[2^{-j}m_{i},2^{-j}(m_{i}+1))$. For a cube $Q$, $\ell(Q)$ stands for the length of cube $Q$. We denote by $\mathcal{D}$ the the family of all dyadic cubes in $\mathbb{R}^{n}$, while $\mathcal{D}_{j}$ is the set of all dyadic cubes with $\ell(Q)=2^{-j},j\in\mathbb{Z}$. Let $x_{Q}$ be the left corner of $Q\in{\mathcal{D}}$. For $\lambda>0$, let $\lambda Q$ be the cube with the same center of $Q$ and the edge length $\lambda\ell(Q)$. For $k\in\mathbb{N}$, let $k_{\operatorname{pa}}Q$ be the $k$-th dyadic parent of $Q$, which is the dyadic cube in ${\mathcal{D}}$ satisfying $Q\subset k_{\operatorname{pa}}Q$ and $\ell(k_{\operatorname{pa}}Q)=2^{k}\ell(Q)$.
2.1 Matrix weights
First, we recall some basic concepts and results from the theory of matrix weights.
For any $d,n\in\mathbb{N}$, denote by $M_{d,n}(\mathbb{C})$ the set of all $d\times n$ complex-valued matrices. $M_{d,d}(\mathbb{C})$ is simply denoted by $M_{d}(\mathbb{C})$. The zero matrix in $M_{d,n}(\mathbb{C})$ (or $M_{d}(\mathbb{C})$) is denoted by $O_{d,n}$ (or $O_{d}$). Denote by $A^{*}$ the conjugate transpose of $A\in M_{d,n}(\mathbb{C})$. A matrix $A\in M_{d}(\mathbb{C})$ is called a Hermitian matrix if $A=A^{*}$ and is called a unitary matrix if $A^{*}A=I_{d}$ where $I_{d}$ is the identity matrix. We denote a diagonal matrix by $\operatorname{diag}(\lambda_{1},\ldots,\lambda_{d})=\operatorname{diag}(\lambda_{i}).$
For a vector $x\in\mathbb{C}^{d}$, let $|x|=(\sum_{i=1}^{d}|x_{i}|^{2})^{1/2}.$ For $1\leq p<\infty$, let $|x|_{p}=(\sum_{i=1}^{d}|x_{i}|^{p})^{1/p}$. For $p=\infty$, let $|x|_{\infty}=\max(x_{1},\ldots,x_{d})$. In the finite dimension space, the norms are equivalent. That is, for $1\leq p<q\leq\infty$,
$|x|_{q}\leq|x|_{p}\leq d^{1/p}|x|_{\infty}\leq d^{1/p}|x|_{q}.$ (2.1)
For $A\in M_{d}(\mathbb{C})$, let
$\left\|A\right\|:=\sup_{\vec{z}\in\mathbb{C}^{d},\left|\vec{z}\right|=1}\left|A\vec{z}\right|.$
We say that a matrix $A\in M_{d}(\mathbb{C})$ is positive definite if, for any $\vec{x}\in\mathbb{C}^{d}\backslash\{\vec{0}\}$, $\vec{z}^{*}A\vec{z}>0.$ And a matrix $A\in M_{d}(\mathbb{C})$ is called nonnegative definite if, for any $\vec{x}\in\mathbb{C}^{d}$, $\vec{z}^{*}A\vec{z}\geq 0.$
From [18, Theorem 4.1.4], any nonnegative definite matrix is always Hermitian. Hence any nonnegative definite matrix is self-adjoint.
Let $A\in M_{d}(\mathbb{C})$ be a positive definite matrix and have eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$. From [18, Theorem 2.5.6(c)], there exists a unitary matrix $U\in M_{d}(\mathbb{C})$ such that
$A=U\operatorname{diag}\,(\lambda_{1},\ldots,\lambda_{d})U^{*}.$ (2.2)
Moreover, by [18, Theorem 4.1.8], we find $\{\lambda_{i}\}_{i=1}^{d}\subset(0,\infty)^{d}$. The following definition is based on these conclusions.
Definition 2.1
Let $A\in M_{d}(\mathbb{C})$ be a positive definite matrix with positive eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$. For any $\alpha\in\mathbb{R}$, define
$A^{\alpha}:=U\operatorname{diag}\left(\lambda_{1}^{\alpha},\ldots,\lambda_{m}^{\alpha}\right)U^{*},$ (2.3)
where $U$ is the same as in (2.2).
Remark 2.2
From [17, p. 408], we obtain that $A^{\alpha}$ is independent of the choices of the order of $\{\lambda_{i}\}_{i=1}^{m}$ and $U$, and hence $A^{\alpha}$ is well defined.
Now, we recall some concepts of matrix wights.
Definition 2.3
A matrix-valued function $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ is called a matrix weight if $W$ satisfies that
(i) for any $x\in{\mathbb{R}^{n}}$, $W(x)$ is nonnegative definite;
(ii) for almost every $x\in{\mathbb{R}^{n}}$, $W(x)$ is invertible;
(iii) the entries of $W$ is locally integrable;
Definition 2.4
Let $p\in(0,\infty)$, $W$ be a matrix weight. Suppose that $E\subset\mathbb{R}^{n}$ is a bounded measurable set satisfying $0<|E|<\infty$. Then the matrix $A_{E}\in M_{d}(\mathbb{C})$ is called a reducing operator of order $p$ for $W$ if $A_{E}$ is positive definite and, for any $\vec{z}\in\mathbb{C}^{d}$,
$\left|A_{E}\vec{z}\right|\approx\left(\frac{1}{|E|}\int_{E}\left|W^{\frac{1}{p}}(x)\vec{z}\right|^{p}\,{\mathrm{d}}x\right)^{\frac{1}{p}},$ (2.4)
where the positive equivalence constants depend only on $d$ and $p$.
Next we recall the concepts of scalar weight class $A_{p}$ (see [12, Definitions 7.1.1, 7.1.3]) and matrix weight class $\mathcal{A}_{p}$ (see [26] for $1<p<\infty$, [9] for $0<p\leq 1$).
Definition 2.5
A weight $\omega$ is a nonnegative locally integrable function on ${\mathbb{R}^{n}}$ such that $0<\omega(x)<\infty$ for almost all $x\in{\mathbb{R}^{n}}$.
A weight $\omega$ is called an $A_{1}$ weight if $\mathcal{M}(\omega)(x)\leq c\omega(x)$ for almost all $x\in{\mathbb{R}^{n}}$ where $\mathcal{M}$ is the Hardy-Littlewood maximal operator.
For $1<p<\infty$, a weight $\omega$ is said to be of class $A_{p}$ if
$\underset{Q\;\operatorname{cubes\;in}{\mathbb{R}^{n}}}{\sup}\left(\frac{1}{|Q|}\int_{Q}\omega(x){\mathrm{d}}x\right)\left(\frac{1}{|Q|}\int_{Q}\omega(x)^{-1/(p-1)}{\mathrm{d}}x\right)^{p-1}<\infty.$
For $1<p<\infty$, a matrix weight $W\in\mathcal{A}_{p}(\mathbb{\mathbb{R}}^{n})$ if and only if
$\sup_{Q}\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|Q|}\int_{Q}\|W^{1/p}(x)W^{-1/p}(y)\|^{p^{\prime}}\mathrm{d}y\right)^{p/p^{\prime}}\mathrm{d}x<\infty,$
where $p^{\prime}=p/(p-1)$ is the conjugate index of $p$, and the supremum is taken over all cubes $Q\subset\mathbb{R}^{n}$.
For $0<p\leq 1$, a matrix weight $W\in\mathcal{A}_{p}(\mathbb{\mathbb{R}}^{n})$ if and only if
[…]
We write $\mathcal{A}_{p}:=\mathcal{A}_{p}(\mathbb{\mathbb{R}}^{n})$ for brevity.
Given any matrix weight $W$ and $0<p<\infty$, there exists (see e.g., [11, Proposition 1.2]) for $p>1$ and [9, p.1237] for $0<p\leq 1)$ a sequence $\{A_{Q}\}_{Q\in\mathcal{D}}$ of positive definite $d\times d$ matrices such that
$c_{1}|A_{Q}\vec{y}|\leq\Big{(}\frac{1}{|Q|}\int_{Q}|W^{1/p}(x)\vec{y}|^{p}\mathrm{d}x\Big{)}^{1/p}\leq c_{2}|A_{Q}\vec{y}|,$
with positive constants $c_{1}$, $c_{2}$ independent of $\vec{y}\in\mathbb{C}^{d}$ and $Q\in\mathcal{D}$. In this case, we call $\{A_{Q}\}_{Q\in\mathcal{D}}$ a sequence of reducing operators of order $p$ for $W$.
Definition 2.6
A matrix weight $W$ is called a doubling matrix weight of order $p>0$ if the scalar measures $w_{\vec{y}}(x)=|W^{1/p}(x)\vec{y}|^{p}$, for $\vec{y}\in\mathbb{C}^{d}$, are uniformly doubling: there exists $c>0$ such that for all cubes $Q\subset\mathbb{R}^{n}$ and all $\vec{y}\in\mathbb{C}^{d}$,
[…]
If $c=2^{\beta}$ is the smallest constant for which this inequality holds, we say that $\beta$ is the doubling exponent of $W$.
From [15, Proposition 2.10], we know that $\beta$ is always not less than $n$.
In [3], Bu, Hytönen, Yang and Yuan introduced the $\mathcal{A}_{p}$-dimension of matrix weights, which will be used in Theorems 4.5 and 4.6.
Definition 2.7
Let $0<p<\infty$, $\tilde{d}\in\mathbb{R}$. A matrix weight $W$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}$, denoted by $W\in\mathbb{D}_{p,\tilde{d}}(\mathbb{R}^{n},\mathbb{C}^{n})$, if there exists a positive constant $C$ such that for any cube $Q\subset\mathbb{R}^{n}$ and any $i\in\mathbb{N}_{0}$,
[…]
or,
[…]
where $11/p+1/p^{\prime}=$ 1.
We denote $\mathbb{D}_{p,\tilde{d}}(\mathbb{R}^{n},\mathbb{C}^{n})$ simply by $\mathbb{D}_{p,\tilde{d}}$.
Lemma 2.8 (Corollary 2.32, [3])
Let $0<p<\infty$, let $W\in\mathcal{A}_{p}$ have the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$, and let $\{A_{Q}\}_{\operatorname{cube}Q}$ be the reducing operator of order $p$ for $W$. (i) If $1<p<\infty$, let $\widetilde{W}:=W^{-1/(p-1)}$ (which belongs to $\mathcal{A}_{p^{\prime}}$) have the $\mathcal{A}_{p^{\prime}}$-dimension $\tilde{\tilde{d}}$. Then there exists a positive constant $C$ such that, for any cubes $Q$ and $R$ of ${\mathbb{R}^{n}}$,
$\|A_{Q}A_{R}^{-1}\|\leq C\max\left(\left[\frac{\ell(R)}{\ell(Q)}\right]^{\tilde{d}/p},\left[\frac{\ell(Q)}{\ell(R)}\right]^{\tilde{\tilde{d}}/p^{\prime}}\right)\left[1+\frac{|x_{Q}-x_{R}|}{\max(\ell(Q),\ell(R))}\right]^{\tilde{d}/p+\tilde{\tilde{d}}/p^{\prime}}.$
(ii) If $0<p\leq 1$, then there exists a positive constant $C$ such that, for any cubes $Q$ and $R$ of ${\mathbb{R}^{n}}$,
$\|A_{Q}A_{R}^{-1}\|\leq C\max\left(\left[\frac{\ell(R)}{\ell(Q)}\right]^{\tilde{d}/p},1\right)\left[1+\frac{|x_{Q}-x_{R}|}{\max(\ell(Q),\ell(R))}\right]^{\tilde{d}/p}.$
2.2 Matrix weighted Bourgain-Morrey spaces
Definition 2.9
Let $0<p<\infty$, $d\in\mathbb{N}$ and $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. The space $L^{p}_{\operatorname{loc}}(W)$ collects all measurable functions $\vec{f}:{\mathbb{R}^{n}}\to\mathbb{C}^{d}$ such that for each compact set $K$,
$\|\vec{f}\chi_{K}\|_{L^{p}(W)}=\left(\int_{K}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}<\infty.$
Let $\varOmega\subset{\mathbb{R}^{n}}$ be an open set. The space $L^{p}(W,\varOmega)$ collects all measurable functions $\vec{f}:\varOmega\to\mathbb{C}^{d}$ such that
$\|\vec{f}\|_{L^{p}(W,\varOmega)}=\left(\int_{\varOmega}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}<\infty.$
Definition 2.10
Let ${\mathcal{D}}=\{Q_{j,k}\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}$ be the standard dyadic system. Let $0<p<t<r<\infty$ or $0<p\leq t<r=\infty$. Let $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. Define $M_{p}^{t,r}(W)$ as the set of all $\vec{f}\in L_{\rm loc}^{p}(W)$ such that
$\|\vec{f}\|_{M_{p}^{t,r}(W)}:=\left\|\left\{W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(y)\vec{f}(y)|^{p}{\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty,$
where $W(Q_{j,k})=\int_{Q_{j,k}}\|W(y)\|{\mathrm{d}}y$.
Let $\{A_{Q}\}_{Q\in{\mathcal{D}}}$ be the reducing operator of order $p$ for $W$. Define $M_{p}^{t,r}(\{A_{Q}\})$ as the set of all $\vec{f}\in L_{\rm loc}^{p}(\{W\})$ such that
$\|\vec{f}\|_{M_{p}^{t,r}(\{A_{Q}\})}$
:=$:=\left\|\left\{\left(\|A_{Q_{j,k}}\|^{p}|Q_{j,k}|\right)^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(y)|^{p}{\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$
Remark 2.11
Let $\{\vec{e}_{i}\}_{i=1}^{d}$ be the standard unit basis for $\mathbb{C}^{d}$. Since $W$ is a matrix weight, then for each $x\in{\mathbb{R}^{n}}$, $W(x)$ is a nonnegative matrix and self-adjoint. Hence we have
$\|W(x)\|$ $=\|W^{2/p}(x)\|^{p/2}\approx[\operatorname{tr}(W^{2/p}(x))]^{p/2}$
$=\left(\sum_{i=1}^{d}|W^{1/p}(x)\vec{e}_{i}|^{2}\right)^{p/2}\approx\sum_{i=1}^{d}|W^{1/p}(x)\vec{e}_{i}|^{p}.$
From this and (2.4), we obtain
$W(Q_{j,k})$ $=\int_{Q_{j,k}}\|W(y)\|{\mathrm{d}}y\approx\int_{Q_{j,k}}\sum_{i=1}^{d}|W^{1/p}(x)\vec{e}_{i}|^{p}{\mathrm{d}}y$
$\approx\sum_{i=1}^{d}|A_{Q_{j,k}}\vec{e}_{i}|^{p}|Q_{j,k}|\approx\|A_{Q_{j,k}}\|^{p}|Q_{j,k}|.$
Thus $M_{p}^{t,r}(W)$ is same with $M_{p}^{t,r}(\{A_{Q}\})$ in meaning of the equivalent quasi-norms.
If $d=1,W\equiv 1$, $M_{p}^{t,r}(W)$ is the classical Bourgain-Morrey space $M_{p}^{t,r}$ in [16]. We define the scalar weighted Bourgain-Morrey space $M_{p}^{t,r}(\omega)$ by
$\|f\|_{M_{p}^{t,r}(\omega)}:=\left\|\left\{(\omega(Q_{j,k}))^{1/t-1/p}\left(\int_{Q_{j,k}}|f(y)|^{p}\omega(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$ (2.6)
That is $M_{p}^{t,r}(\omega)$ is the case $d=1$ of matrix weighted Bourgain-Morrey space $M_{p}^{t,r}(W)$.
The following Lemma is proved in [22, 28].
Lemma 2.12
Let $1\leq p<\infty$ and $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. Let $\varOmega\subset{\mathbb{R}^{n}}$ be an open set. Then matrix weighted Lebesgue space $L^{p}(W,\varOmega)$ is a Banach space.
Theorem 2.13
Let $1\leq p<\infty$ and $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. Then the space $L^{p}_{\operatorname{loc}}(W)$ is complete.
  • Proof
    Let $\{\vec{f}_{k}\}_{k=1}^{\infty}$ be a Cauchy sequence in $L^{p}_{\operatorname{loc}}(W)$. That is, for any compact set $K$, any $\varepsilon>0$, there exists $N>0$ such that if $j,k>N$, $\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\varepsilon.$ Then for any compact set $K$, $\{\vec{f}_{k}\chi_{K}\}_{k=1}^{\infty}$ is a Cauchy sequence in $L^{p}(W)$. Since $L^{p}(W)$ is complete (Lemma 2.12), there exists $\vec{f}\chi_{K}$ in $L^{p}(W)$ such that $\vec{f}_{k}\chi_{K}\to\vec{f}\chi_{K}$. By the Fatou lemma, we have
    $\|\vec{f}\chi_{K}\|_{L^{p}(W)}\leq\liminf_{k\to\infty}\|\vec{f}_{k}\chi_{K}\|_{L^{p}(W)}<\infty.$
    By the dominated convergence theorem, for $j,k>N$, we obtain
    $\|(\vec{f}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}=\lim_{j\to\infty}\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\varepsilon.$
    Hence $\vec{f}_{k}\to\vec{f}$ in $L^{p}_{\operatorname{loc}}(W)$ and the space $L^{p}_{\operatorname{loc}}(W)$ is complete. $\Box$
In what follows, the symbol $\hookrightarrow$ always stands for continuous embedding.
Proposition 1
Let $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight.
(i) If $0<p<t<r_{1}<r_{2}\leq\infty$, then
$M_{p}^{t,r_{1}}(W)\hookrightarrow M_{p}^{t,r_{2}}(W).$
(ii) If $0<p_{1}<p_{2}<t<r<\infty$ or $0<p_{1}<p_{2}\leq t<r=\infty$, then
$M_{p_{2}}^{t,r}(W)\hookrightarrow M_{p_{1}}^{t,r}(W).$
(iii) If $0<p\leq t<r=\infty$, then
$L^{t}(W)\hookrightarrow M_{p}^{t,r}(W)\hookrightarrow L_{\operatorname{loc}}^{p}(W).$
  • Proof
    (i) It comes from $\ell^{r_{1}}\hookrightarrow\ell^{r_{2}}$ since $0<r_{1}<r_{2}\leq\infty$.
    (ii) It comes from the Hölder inequality. Indeed, for each $Q\in{\mathcal{D}}$,
    $\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{1}}(y)\vec{f}(y)|^{p_{1}}{\mathrm{d}}y\right)^{1/p_{1}}$
    $=\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{1}-1/p_{2}}(y)W^{1/p_{2}}(y)\vec{f}(y)|^{p_{1}}{\mathrm{d}}y\right)^{1/p_{1}}$
    $\leq\left(\frac{1}{W(Q)}\int_{Q}\|W^{1/p_{1}-1/p_{2}}(y)\|^{p_{1}}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{1}}{\mathrm{d}}y\right)^{1/p_{1}}$
    $\leq\left(\frac{1}{W(Q)}\int_{Q}\|W^{1/p_{1}-1/p_{2}}(y)\|^{p_{1}p_{2}/(p_{2}-p_{1})}{\mathrm{d}}y\right)^{1/p_{1}-1/p_{2}}$
    ×$\quad\times\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{2}}{\mathrm{d}}y\right)^{1/p_{2}}$
    $=\left(\frac{1}{W(Q)}\int_{Q}\|W(y)\|{\mathrm{d}}y\right)^{1/p_{1}-1/p_{2}}\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{2}}{\mathrm{d}}y\right)^{1/p_{2}}$
    $=\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{2}}{\mathrm{d}}y\right)^{1/p_{2}}.$
    Hence,
    $\|\vec{f}\|_{M_{p_{1}}^{t,r}(W)}\leq\|\vec{f}\|_{M_{p_{2}}^{t,r}(W)}.$
    Thus we prove (ii).
    (iii) The first embedding comes from the fact that
    $L^{t}(W)=M_{t}^{t,\infty}(W)\hookrightarrow M_{p}^{t,\infty}(W).$
    For any compact set $K\subset{\mathbb{R}^{n}}$, there exist at most $2^{n}$ dyadic cubes $Q_{j}$ such that $K\subset\cup_{j=1}^{2n}Q_{j}$. Hence
    $\|\vec{f}\chi_{K}\|_{L^{p}(W)}$ $\leq\sum_{j=1}^{2^{n}}\frac{|Q_{j}|^{1/t-1/p}}{|Q_{j}|^{1/t-1/p}}\|\vec{f}\chi_{Q_{j}}\|_{L^{p}(W)}$
    $\leq\|\vec{f}\|_{M_{P}^{t,\infty}(W)}\sum_{j=1}^{2^{n}}\frac{1}{|Q_{j}|^{1/t-1/p}}<\infty.$
    This shows that $M_{p}^{t,r}(W)\hookrightarrow L_{\operatorname{loc}}^{p}(W).$
    $\Box$
Proposition 2
The matrix weighted Bourgain-Morrey space $M_{p}^{t,r}(W)$ is a Banach space when $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$.
  • Proof
    The argument are standard, see, for example, [27, Theorem 2.4]. We only show that $M_{p}^{t,r}(W)$ is complete since others are simple.
Let $\{\vec{f}_{j}\}_{j=1}^{\infty}$ be a Cauchy sequence in $M_{p}^{t,r}(W)$. By Proposition 1, we have that $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is also a Cauchy sequence in $L^{p}_{\operatorname{loc}}(W)$. Since $L^{p}_{\operatorname{loc}}(W)$ is complete, there exists a vector function $\vec{f}$ in $L^{p}_{\operatorname{loc}}(W)$ such that $\vec{f}_{j}\to\vec{f}$. By the Fatou lemma,
$\|\vec{f}\|_{M_{p}^{t,r}(W)}=\|\underset{j\to\infty}{\lim\inf}\vec{f}_{j}\|_{M_{p}^{t,r}(W)}\leq\underset{j\to\infty}{\lim\inf}\|\vec{f}_{j}\|_{M_{p}^{t,r}(W)}<\infty.$
Thus $\vec{f}\in M_{p}^{t,r}(W)$. Therefore,
$\underset{j\to\infty}{\lim\sup}\|\vec{f}-\vec{f}_{j}\|_{M_{p}^{t,r}(W)}\leq\underset{j\to\infty}{\lim\sup}\left(\underset{k\to\infty}{\lim\inf}\|\vec{f}_{k}-\vec{f}_{j}\|_{M_{p}^{t,r}(W)}\right)=0.$
Thus, the sequence $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is convergent to $\vec{f}$ in $M_{p}^{t,r}(W)$. $\Box$
Recall that a matrix weight $W$ is almost everywhere invertible. Hereafter, we define $w(x):=\|W^{-1}(x)\|^{-1}$ when a matrix weight $W$ is invertible at $x\in{\mathbb{R}^{n}}$.
Lemma 2.14 (Proposition 3.2,[8])
For a matrix weight $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$, we have $0<w(x)\leq\|W(x)\|<\infty$ for a.e. $x\in{\mathbb{R}^{n}}$. Furthermore, $W$ satisfies a two weight, degenerate ellipticity condition: for $\xi\in\mathbb{R}^{d}$,
$w(x)|\xi|^{p}\leq|W^{1/p}(x)\xi|^{p}\leq\|W(x)\||\xi|^{p}.$ (2.7)
Remark 2.15
In [8, Proposition 3.2], it is assumed that $\xi\in\mathbb{R}^{d}$ but it also works for $\xi\in\mathbb{C}^{d}$.
Proposition 3
Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$. Let $W$ be a matrix weight and $\vec{f},\vec{g}\in M_{p}^{t,r}(W)$. Then $\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}=0$ if and only if $\vec{f}(x)=\vec{g}(x)$ a.e..
  • Proof
    Clearly, if $\vec{f}(x)=\vec{g}(x)$ a.e., then $\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}=0$. Then we apply Lemma 2.14 to prove the converse. By the degenerate ellipticity condition (2.7), we have
    00 0 $=\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}$
    $\geq\left\|\left\{W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|\vec{f}(y)-\vec{g}(y)|^{p}w(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}.$
    Since $w(y)>0$ a.e., it follows that $\vec{f}(y)-\vec{g}(y)=0$ a.e.. $\Box$
3 Sufficient conditions for precompact sets
Definition 3.1
(i) Suppose that $(X,\rho)$ is a metric space. Let $A$ be a set of $X$ and let $\varepsilon>0$. A set $E\subset X$ is called an $\varepsilon$-net for $A$ if every point $a\in A$, there exists a point $e\in E$ such that $\rho(a,e)<\varepsilon$.
(ii) A set $A$ is called totally bounded if, for each $\varepsilon>0$, it possesses a finite $\varepsilon$-net.
(iii) A subset in a topological space is precompact if its closure is compact.
It is well known that in a complete metric space, a set is precompact if and only if it is totally bounded.
Theorem 3.2
Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$. Let $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. A subset $\mathcal{F}\subset M_{p}^{t,r}(W)$ is totally bounded if the following conditions are valid:
(i) $\mathcal{F}$ uniformly vanishes at infinity, that is,
$\lim_{R\to\infty}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}=0;$
(ii) $\mathcal{F}$ is equicontinuous, that is,
$\lim_{b\to 0}\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in B(0,b)}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}=0.$ (3.1)
Here and what follows, $\tau_{y}$ denotes the translation operator: $\tau_{y}\vec{f}(x):=\vec{f}(x-y)$.
Remark 3.3
In this Remark, we will prove that the conditions (i) and (ii) in Theorem 3.2 together imply that the set $\mathcal{F}$ is bounded. Indeed, choose $b>0$ such that for all $h\in{\mathbb{R}^{n}},|h|\leq b$, all $\vec{f}\in\mathcal{F}$,
$\|\vec{f}-\tau_{y}\vec{f}\|_{M^{t,r}_{p}(W)}\leq 1.$
Choose $R>0$ such that for all $\vec{f}\in\mathcal{F}$,
$\|\vec{f}\chi_{B^{c}(0,R)}\|_{M^{t,r}_{p}(W)}\leq 1.$
Fix $h$ with $|h|=b$. Then for all $\vec{f}\in\mathcal{F}$, $x\in{\mathbb{R}^{n}}$, we have
$\|\vec{f}\chi_{B(x,R)}\|_{M^{t,r}_{p}(W)}$ $\leq\|(\vec{f}-\tau_{h}\vec{f})\chi_{B(x,R)}\|_{M^{t,r}_{p}(W)}+\|\tau_{h}\vec{f}\chi_{B(x,R)}\|_{M^{t,r}_{p}(W)}$
$=\|(\vec{f}-\tau_{h}\vec{f})\chi_{B(x,R)}\|_{M^{t,r}_{p}(W)}+\|\vec{f}\chi_{B(x+h,R)}\|_{M^{t,r}_{p}(W)}$
$\leq 1+\|\vec{f}\chi_{B(x+h,R)}\|_{M^{t,r}_{p}(W)}.$
Hence, by induction,
$\|\vec{f}\chi_{B(0,R)}\|_{M^{t,r}_{p}(W)}\leq N+\|\vec{f}\chi_{B(Nh,R)}\|_{M^{t,r}_{p}(W)}.$
Now choose $N\geq 1$ such that $Nb=N|h|>2R$. Then $B(Nh,R)\subset B^{c}(0,R)$. Hence
$\|\vec{f}\|_{M^{t,r}_{p}(W)}$ $\leq\|\vec{f}\chi_{B(0,R)}\|_{M^{t,r}_{p}(W)}+\|\vec{f}\chi_{B^{c}(0,R)}\|_{M^{t,r}_{p}(W)}$
$\leq N+\|\vec{f}\chi_{B(Nh,R)}\|_{M^{t,r}_{p}(W)}+\|\vec{f}\chi_{B^{c}(0,R)}\|_{M^{t,r}_{p}(W)}\leq N+2.$
This proves that $\mathcal{F}$ is bounded.
  • Proof of Theorem 3.2.
    Assume that $\mathcal{F}\subset M_{p}^{t,r}(W)$ satisfies (i) and (ii). Given $\varepsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\varepsilon$-net of $\mathcal{F}$. Denote by $R_{i}:=[-2^{i},2^{i})^{n}$ for $i\in\mathbb{Z}$. Then from condition (i), there exist a positive integer $m$ large enough such that
    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (3.2)
    Moreover, by condition (ii), there exists a negative integer $a$ such that
    $\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R_{a}}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (3.3)
    There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\cup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$.
    For any $\vec{f}\in\mathcal{F}$ and $x\in{\mathbb{R}^{n}}$, let
    $\varPhi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
    Then for each fixed $x\in{\mathbb{R}^{n}}$, we have
    $\left|W(x)\left(\vec{f}(x)-\vec{f}_{Q_{j}}\right)\right|\chi_{Q_{j}}(x)$ $=\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}W(x)\left(\vec{f}(x)-\vec{f}(y)\right){\mathrm{d}}y\right|\chi_{Q_{j}}(x)$ (3.4)
    $\leq\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x).$ (3.5)
    We split $\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}$ into three parts:
    $\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    $\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    +$\quad+\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    +$\quad+\left\{\sum_{k<-m}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    $=:S_{1}+S_{2}+S_{3}.$
    We first estimate $S_{1}$. By (3.4), Jensen’s inequality ($p\geq 1$), the Fubini theorem, we have
    $S_{1}$$\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.6)
    ×$\quad\times\left.\left(\int_{Q}\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.7)
    $\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.8)
    ×$\quad\times\left.\left(\frac{1}{|Q_{j}|}\int_{Q_{j}}\int_{Q}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$ (3.9)
    =$=\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.10)
    ×$\quad\times\left.\left(\frac{1}{|Q_{j}|}\int_{Q}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.11)
    $=2^{-an/p}\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.12)
    ×$\quad\times\left.\left(\int_{Q}\int_{x-Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.13)
    where $x-Q_{j}:=\{x-y:y\in Q_{j}\}$. Since $x\in Q\subset Q_{j}$ for some $j\in\{1,2,\ldots,N\}$, we have $x-Q_{j}\subset R_{a}$. Hence by (3.3), we obtain
[…]
    ×$\quad\left.\times\left(\int_{Q}\left|W(x)(\vec{f}(x)-\vec{f}(x-y))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    $\leq\varepsilon.$
    As for $S_{2}$, for each $Q\in{\mathcal{D}}_{k}$ where $k=-m,-m+1,\ldots,-a$, there are $2^{(-k-a)n}$ cubes $Q_{j}\in{\mathcal{D}}_{-a}$ such that $\cup Q_{j}=Q$. Denote by $Q_{j_{\ell}}$, $\ell=1,2,\ldots,2^{(-k-a)n}$ these cubes. Then by (3.4), Jensen’s inequality ($p\geq 1$), the Fubini theorem, we obtain
    $S_{2}$ =$=\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.14)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j_{\ell}}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.15)
    $\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.16)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.17)
    $\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.18)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$ (3.19)
    $=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.20)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.21)
    $=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.22)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{x-Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.23)
    where $x-Q_{j_{\ell}}:=\{x-y:y\in Q_{j_{\ell}}\}$. Note that $x-Q_{j_{\ell}}\subset R_{a}$ when $x\in Q_{j_{\ell}}$. By (3.3), we have
[…]
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    $\leq\varepsilon.$
    As for $S_{3}$, for each set ${\mathcal{D}}_{k}$, there are only $2^{n}$ cubes $Q\in{\mathcal{D}}_{k}$ such that $Q\cap R_{m}\neq\emptyset$. We denote by $Q_{\ell}$, $\ell=1,2,\ldots,2^{n}$ these cubes. And for each cube $Q_{\ell}$, there are $2^{(m-a)n}$ cubes $Q_{j}\in\mathcal{D}_{-a}$ such that $\cup_{j}Q_{j}=Q_{\ell}$. Then
    $S_{3}$ =$=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\int_{Q_{\ell}}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    =$=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\sum_{j=1}^{2^{(m-a)n}}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}.$
    Similarly as $S_{2}$, from (3.4), Jensen’s inequality ($p\geq 1$), the Fubini theorem, and (3.3), we get that $S_{3}\leq\varepsilon$. Together with the estimates of $S_{1}$, $S_{2}$, $S_{3}$, we obtain
    $\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}\leq 3\varepsilon.$ (3.24)
    Note that
    $\|\vec{f}-\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}$ $\leq\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$
    $=\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}.$
    Thus via (3.2) and (3.24), we have
    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq 4\varepsilon.$ (3.25)
    From (3.25), it suffices to show that $\varPhi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$.
    From Remark 3.3, we have
    $\sup_{\vec{f}\in\mathcal{F}}\|\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq\sup_{\vec{f}\in\mathcal{F}}\|\varPhi(\vec{f})-\vec{f}\|_{M_{p}^{t,r}(W)}+\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
    Thus, for any $\vec{f}\in\mathcal{F}$,
    $|W^{1/p}(x)\varPhi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (3.26)
    Since $W$ is a matrix weight, by (i), (ii) of Definition 2.3, we have for almost everywhere $x\in{\mathbb{R}^{n}}$, $W(x)$ is a positive definite matrix. Therefore, we obtain
    $|\varPhi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (3.27)
    which implies $|\vec{f}_{Q_{j}}|<\infty$, $j=1,2,\ldots,N$. From this and the entries of $W$ is locally integrable, we see that $\varPhi$ is a map from $\mathcal{F}$ to $\mathcal{B}$, a finite dimensional Banach subspace of $M_{p}^{t,r}(W)$. Note that $\varPhi(\mathcal{F})\subset\mathcal{B}$ is bounded, and hence is totally bounded. The proof of Theorem 3.2 is complete. $\Box$
Then we give an application in degenerate Bourgain-Morrey-Sobolev spaces with matrix weights.
Definition 3.4
Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$. Let $W:{\mathbb{R}^{n}}\to M_{n}(\mathbb{C})$ be a matrix weight and set the scalar weight $\omega:=\|W\|$. We define the degenerate Bourgain-Morrey-Sobolev spaces $\mathcal{W}^{1,p,t,r}(W)$ by the set of all Lebesgue measurable functions on ${\mathbb{R}^{n}}$ such that
$\|f\|_{\mathcal{W}^{1,p,t,r}(W)}:=\|f\|_{M^{t,r}_{p}(\omega)}+\|\nabla f\|_{M^{t,r}_{p}(W)}<\infty$
where $\nabla f=(\partial_{x_{1}}f,\ldots,\partial_{x_{n}}f)^{T}$ is the gradient of $f$ and
$\|f\|_{M^{t,r}_{p}(\omega)}:=\left\|\left\{\omega(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|f(y)|^{p}\omega(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}},$
is the norm of the scalar weighted Bourgain-Morrey space in (2.6).
Lemma 3.5 (p. 262, [33])
A set $\mathcal{F}$ in metric space $X$ is totally bounded if and only if it is Cauchy-precompact, that is, every sequence has a Cauchy subsequence.
Corollary 3.6
Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$. Let $W:{\mathbb{R}^{n}}\to M_{n}(\mathbb{C})$ be a matrix weight. A subset $\mathcal{F}\subset\mathcal{W}^{1,p,t,r}(W)$ is totally bounded if the following conditions are valid:
(i) $\mathcal{F}$ uniformly vanishes at infinity, that is,
$\lim_{R\to\infty}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\chi_{B^{c}(0,R)}\|_{\mathcal{W}^{1,p,t,r}(W)}=0;$
(ii) $\mathcal{F}$ is equicontinuous, that is,
$\lim_{a\to 0}\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in B(0,a)}\|\vec{f}-\tau_{y}\vec{f}\|_{\mathcal{W}^{1,p,t,r}(W)}=0.$
  • Proof
    Note that $\mathcal{F}\subset\mathcal{W}^{1,p,t,r}(W)$ satisfies conditions (i)-(ii) if and only if $\mathcal{F}\subset M^{t,r}_{p}(\omega)$ satisfies conditions (i)-(ii) of Theorem 3.2 ($d=1$) and :=$\nabla\mathcal{F}:=\{\nabla f:f\in\mathcal{F}\}\subset M^{t,r}_{p}(W)$ satisfies (i)-(ii) of Theorem 3.2 ($d=n$). Hence by Theorem 3.2, we obtain that both $\mathcal{F}\subset M^{t,r}_{p}(\omega)$ and $\nabla\mathcal{F}\subset M^{t,r}_{p}(W)$ are totally bounded. Then Corollary 3.6 follows from Lemma 3.5. $\Box$
4 Dyadic average operator conditions for precompact sets
In this section, we give a criterion for precompactness by dyadic average operator. Let $\mathcal{F}$ be a subset of $M_{p}^{t,r}(W)$. $\mathcal{F}$ is equicontinuous by means of the dyadic average operator if
$\lim_{a\to-\infty,a\in\mathbb{Z}}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=0,$ (4.1)
where
$E_{d,a}\vec{f}(x):=\sum_{Q\in{\mathcal{D}}_{-a}}\frac{\chi_{Q}(x)}{|Q|}\int_{Q}\vec{f}(y){\mathrm{d}}y.$ (4.2)
Recall that $x_{Q}$ is the left corner of $Q\in{\mathcal{D}}$. Then for $x\in Q\in\mathcal{D}_{-a}$
$E_{d,a}\vec{f}(x)=\frac{\chi_{Q_{-a,0}}(x-x_{Q})}{|Q_{-a,0}|}\int_{Q_{-a,0}}\vec{f}(y-x_{Q}){\mathrm{d}}y.$
Then we will prove that (3.1) is stronger than (4.1). Indeed, suppose that condition (3.1) holds. For any $\varepsilon>0$, there exists a cube $R$ with center 00 and side length $02^{a+1}>$ 0 such that
$\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (4.3)
If $x,y\in Q\in{\mathcal{D}}_{-a}$, then $x-y\in R.$ By Jensen’s inequality ($p\geq 1$), Hölder’s inequality ($r/p\geq 1$), (4.3), we obtain
$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|\frac{1}{|Q_{-a,0}|}\right.\right.\right.$
×$\quad\times\left.\left.\left.\int_{Q_{-a,0}}W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\right.\right.$
×$\quad\times\left.\left.\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right)^{1/r}$
$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\right.$
×$\quad\times\left.\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}{\mathrm{d}}y\right)^{1/r}$
$\leq\frac{1}{|Q_{-a,0}|^{1/r}}|Q_{-a,0}|^{1/r}\sup_{y\in R}\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\right.$
×$\quad\times\left.\left(\int_{Q}\left|W^{1/p}(x)(\vec{f}(x)-\vec{f}(x-y))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
$\leq\varepsilon,$
modified when $r=\infty$. Hence we prove that (3.1) is stronger than (4.1).
Replacing the translation operator by the average operator, we have the following result.
Theorem 4.1
Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$. Let $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. A subset $\mathcal{F}\subset M_{p}^{t,r}(W)$ is totally bounded if the following conditions are valid:
(i) $\mathcal{F}$ is bounded, that is,
$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty;$
(ii) $\mathcal{F}$ uniformly vanishes at infinity, that is,
$\lim_{R\to\infty}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}=0;$
(iii) $\mathcal{F}$ is equicontinuous by means of the dyadic average operator, that is, for any
$\lim_{a\to-\infty,a\in\mathbb{Z}}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=0.$
where $E_{d,a}$ is same as (4.2).
  • Proof
    Assume that $\mathcal{F}\subset M_{p}^{t,r}(W)$ satisfies (i)-(iii). Given $\varepsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\varepsilon$-net of $\mathcal{F}$. Denote by $R_{i}:=[-2^{i},2^{i})^{n}$ for $i\in\mathbb{Z}$. Then from condition (ii), there exists a positive integer $m$ large enough such that
    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (4.4)
    Moreover, by condition (iii), there exists an integer $a<0$ such that
    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (4.5)
    There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\cup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$.
    For any $\vec{f}\in\mathcal{F}$ and $x\in{\mathbb{R}^{n}}$, let
    $\varPhi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
    Note that for $x\in R_{m}$,
    $\varPhi(\vec{f})(x)=E_{d,a}\vec{f}(x).$
    Hence via (4.4) and (4.5), we have
    $\|\vec{f}-\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}$ $\leq\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$ (4.6)
    $=\|(\vec{f}-E_{d,a}\vec{f})\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$ (4.7)
    $\leq\|(\vec{f}-E_{d,a}\vec{f})\|_{M_{p}^{t,r}(W)}+\varepsilon$ (4.8)
    $\leq 2\varepsilon.$ (4.9)
    From (4.9), it suffices to show that $\varPhi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$. And we have proved that $\varPhi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$ in Theorem 3.2. Thus the proof of Theorem 4.1 is complete. $\Box$
Lemma 4.2 (Lemma 4.5, [8])
Let $1\leq p<\infty$ and $W\in\mathcal{A}_{p}$. Then $\|W\|$ and $\|W^{-1}\|^{-1}$ are scalar $A_{p}$ weights.
The following is a vector-valued extension of the Lebesgue differentiation theorem on matrix weighted spaces $M^{t,r}_{p}(W)$. For any cube $Q$, $\vec{f}\in L^{1}_{\operatorname{loc}}({\mathbb{R}^{n}})$, define
$E_{Q}(\vec{f})(x)=\frac{1}{|Q|}\int_{Q}\vec{f}(y){\mathrm{d}}y.$
Theorem 4.3
Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$. If $W\in\mathcal{A}_{p}$, then for any $\vec{f}\in M_{p}^{t,r}(W)$,
$\lim_{\ell(Q)\to 0}\left|E_{Q}\vec{f}(x)-\vec{f}(x)\right|=0\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$
  • Proof
    First, for any $\vec{f}=(f_{1},f_{2},\ldots,f_{d})^{T}\in M_{p}^{t,r}(W)$, it suffices to show that $f_{i}\in L^{1}_{\operatorname{loc}}({\mathbb{R}^{n}})$ for each $i=1,2,\ldots,d.$ Since $W\in\mathcal{A}_{p}$, then by (2.3)
    $|\vec{f}|^{p}=|W^{-1/p}W^{1/p}\vec{f}|^{p}\leq\|W^{-1/p}\|^{p}|W^{1/p}\vec{f}|^{p}=\|W^{-1}\||W^{1/p}\vec{f}|^{p}.$
    It follows that
    $|\vec{f}|^{p}\|W^{-1}\|^{-1}\leq|W^{1/p}\vec{f}|^{p}.$
    From Lemma 4.2, we conclude that $\|W^{-1}\|^{-1}$ is a scalar $A_{p}$ weight, hence $|\vec{f}|\in L^{1}_{\operatorname{loc}}({\mathbb{R}^{n}})$. Indeed, since any compact set $K\subset{\mathbb{R}^{n}}$ can be contained in a finite set of dyadic cubes, it suffices to show that for any $Q\in{\mathcal{D}}$, $|\vec{f}|\in L^{1}(Q)$. By Hölder’s inequality, we have
    $\int_{Q}|\vec{f}|{\mathrm{d}}x$ $=\int_{Q}|\vec{f}|\|W^{-1}\|^{-1/p}\|W^{-1}\|^{1/p}{\mathrm{d}}x$
    $\leq\left(\int_{Q}|\vec{f}|^{p}\|W^{-1}\|^{-1}{\mathrm{d}}x\right)^{1/p}\left(\int_{Q}\|W^{-1}\|^{p^{\prime}/p}{\mathrm{d}}x\right)^{1/p^{\prime}}$
    $\leq C_{Q,W,p}\left(\int_{Q}|W^{1/p}\vec{f}|^{p}{\mathrm{d}}x\right)^{1/p}$
    $\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
    Hence, $f_{i}\in L^{1}_{\operatorname{loc}}({\mathbb{R}^{n}})$ for each $i=1,2,\ldots,d$. By the classical Lebesgue differentiation theorem, for each $1\leq i\leq d$, we have
    $\lim_{\ell(Q)\to 0}\left|E_{Q}f_{i}(x)-f_{i}(x)\right|=0\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$
    Theorem 4.3 comes from the fact that for any $x\in{\mathbb{R}^{n}}$,
    $\left|E_{Q}\vec{f}(x)-\vec{f}(x)\right|\leq d^{1/2}\max_{i}\left|E_{Q}f_{i}(x)-f_{i}(x)\right|.$
Next we need the boundedness of dyadic average operator. Given any collection $\mathcal{Q}$ of pairwise disjoint cubes $Q\subset{\mathbb{R}^{n}}$, define the averaging operator $A_{\mathcal{Q}}$ by
$A_{\mathcal{Q}}\vec{f}(x)=\sum_{Q\in\mathcal{Q}}\frac{\chi_{Q}(x)}{|Q|}\int_{Q}\vec{f}(y){\mathrm{d}}y.$
Lemma 4.4 (Proposition 4.7,[8])
Let $1\leq p<\infty$. Let $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. Then $W\in\mathcal{A}_{p}$ if and only it satisfies
$\|A_{\mathcal{Q}}\vec{f}\|_{L^{p}(W)}\leq C_{n,d,p,W}\|\vec{f}\|_{L^{p}(W)}.$
Theorem 4.5
Let $W\in\mathcal{A}_{p}$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$ with the doubling exponent $\beta$.
(i) Let $1\leq p<t<r<\infty$ If $(\tilde{d}-n+\beta)r/p+n-\beta r/t<0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
(ii) Let $1\leq p\leq t<r=\infty$ If $\tilde{d}/p-n/p-\beta(1/t-1/p)\leq 0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
  • Proof
    (i) For a fixed $a\in\mathbb{Z}$, we have
    $\|E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}$ $\leq\left(\sum_{j\leq-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
    $+\left(\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
    $=:S_{1}+S_{2}.$
    By Lemma 4.4, we obtain
    $S_{1}$ $\leq C_{n,d,p,W}\left(\sum_{j\leq-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
    $\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}.$
    When $j>-a$, we denote by $(j+a)_{\operatorname{pa}}Q\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q\in{\mathcal{D}}_{j}$. Let $\{A_{Q}\}_{Q\in{\mathcal{D}}}$ be the reducing operator of order $p$ for $W$. By Lemmas 2.8, we have
    $\|A_{Q}A_{(j+a)_{\operatorname{pa}}Q}^{-1}\|\leq c2^{(j+a)\tilde{d}/p}.$ (4.10)
    In $S_{2}$, using (2.5) and (4.10) , we have
    $S_{2}^{r}$ $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
    $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}$
    ×$\quad\times\left(\frac{|Q|}{|Q|}\int_{Q}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}$
    $\approx\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}\left(\left|A_{Q}\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|\right)^{r}$
    $\lesssim\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}$
    ×$\quad\times\left(2^{(j+a)\tilde{d}/p}\left|A_{(j+a)_{\operatorname{pa}}Q}\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|\right)^{r}$
    $\approx\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}2^{(j+a)\tilde{d}r/p}$
    ×$\quad\times\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y|^{p}{\mathrm{d}}x\right)^{r/p}$
    $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}$
    ×$\quad\times\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
    $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}$
    ×$\quad\times\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}.$
    Remark that given $S\in{\mathcal{D}}_{-a}$, there are $2^{(j+a)n}$ cubes $R$ such that $(j+a)R=S$. Hence, by Lemma 4.4 and $(\tilde{d}-n+\beta)r/p+n-\beta r/t<0$, we have
    $\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}$
    ×$\times\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
    $=\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}$
    ×$\quad\times\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
    $\leq\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}2^{(j+a)n}2^{-(j+a)\beta(r/t-r/p)}$
    ×$\quad\times\sum_{S\in\mathcal{D}_{-a}}W(S)^{r/t-r/p}\left(\int_{S}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
    $\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}^{r},$
    where in the second step we use the following fact. Since $W((j+a)_{\operatorname{pa}}Q)\leq 2^{(j+a)\beta}W(Q)$ and $01/t-1/p<$ 0, we have
    $W(Q)^{1/t-1/p}\leq 2^{-(j+a)\beta(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p}.$
(ii) Fix $Q_{j,k}\in{\mathcal{D}}$. If $j\leq-a$, then by Lemma 4.4, we have
$W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}$
$\lesssim W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}.$
If $j>-a$, we denote by $(j+a)_{\operatorname{pa}}Q_{j,k}\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q_{j,k}\in{\mathcal{D}}_{j}$.
$W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}$
$=W(Q_{j,k})^{1/t-1/p}\left(\frac{|Q_{j,k}|}{|Q_{j,k}|}\int_{Q_{j,k}}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{1/p}$
$\approx W(Q_{j,k})^{1/t-1/p}|Q_{j,k}|^{1/p}\left|A_{Q_{j,k}}\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\vec{f}(y){\mathrm{d}}y\right|$
$\lesssim W(Q_{j,k})^{1/t-1/p}|Q_{j,k}|^{1/p}2^{(j+a)\tilde{d}/p}\left|A_{(j+a)_{\operatorname{pa}}Q_{j,k}}\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\vec{f}(y){\mathrm{d}}y\right|$
$\approx W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\right.$
×$\quad\times\left.\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{1/p}.$
Then by Lemma 4.4 and $\tilde{d}/p-n/p-\beta(1/t-1/p)\leq 0$, we obtain
$W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}$
$\lesssim W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}$
×$\quad\times\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
$=W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}2^{-an/p}\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
$\lesssim 2^{(j+a)(\tilde{d}/p-n/p)}2^{-(j+a)\beta(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}$
×$\quad\times\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
$\lesssim W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
where in the third step we use the following fact. Since $W((j+a)_{\operatorname{pa}}Q)\leq 2^{(j+a)\beta}W(Q)$ and $1/t-1/p\leq 0$, we have
$W(Q)^{1/t-1/p}\leq 2^{-(j+a)\beta(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p}.$
Finally, take the supremum over the cubes $Q_{j,k}\in D$, we obtain
$\|E_{d,a}\vec{f}\|_{M_{p}^{t,\infty}(W)}\lesssim\|\vec{f}\|_{M_{p}^{t,\infty}(W)}.$
Finally, we have the following Kolmogorov-Riesz compactness theorem for matrix weighted Bourgain-Morrey spaces.
Theorem 4.6
Let $1\leq p<t<r<\infty$ Let $W\in\mathcal{A}_{p}$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$ with the doubling exponent $\beta$. Let $(\tilde{d}-n+\beta)r/p+n-\beta r/t<0$. A subset $\mathcal{F}$ of $M_{p}^{t,r}(W)$ is totally bounded if and only if the following conditions hold:
(i) $\mathcal{F}$ is bounded, that is,
$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty;$
(ii) $\mathcal{F}$ uniformly vanishes at infinity, that is,
$\lim_{R\to\infty}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}=0;$
(iii) $\mathcal{F}$ is equicontinuous by means of the dyadic average operator, that is, for any
$\lim_{a\to-\infty,a\in\mathbb{Z}}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=0.$
where $E_{d,a}$ is same as (4.2).
  • Proof
    The sufficiency is due to Theorem 4.1. Now we prove the necessity. Assume that $\mathcal{F}$ of $M_{p}^{t,r}(W)$ is totally bounded. For any given $\varepsilon>0$, there exists $\{\vec{f}_{k}\}_{k=1}^{N_{0}}\subset\mathcal{F}$ such that $\{\vec{f}_{k}\}_{k=1}^{N_{0}}$ is an $\varepsilon$-net of $\mathcal{F}$, that is, for any $\vec{f}\in\mathcal{F}$, there exists $\vec{f}_{k},k\in\{1,2,\ldots,N_{0}\}$ such that $\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}<\varepsilon$.
    Clearly, (i) is true.
    For any $\varepsilon>0$ and $k\in\{1,2,\ldots,N_{0}\}$, since $r<\infty$, there exists $R_{k}>0$ such that
    $\|\vec{f}_{k}\chi_{B^{c}(0,R_{k})}\|_{M_{p}^{t,r}(W)}<\varepsilon.$
    Taking $R=\max_{k\in{1,\ldots,N_{0}}}\{R_{k}\}$, we have $\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ Then for any $\vec{f}\in\mathcal{F}$,
    $\|\vec{f}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}\leq\|(\vec{f}-\vec{f}_{k})\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<2\varepsilon.$
    Thus we prove that
    $\lim_{R\to\infty}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}=0.$
    Therefore, we prove that $\mathcal{F}$ satisfies condition (ii).
As for (iii), for each $1\leq k\leq N_{0}$, by Theorem 4.5, we have for each $a\in\mathbb{Z}$,
$\|\vec{f}_{k}-E_{d,a}\vec{f}_{k}\|_{M_{p}^{t,r}(W)}\leq(1+c)\|\vec{f}_{k}\|_{M_{p}^{t,r}(W)}.$
Thus using the dominated convergence theorem to obtain that there exists $a_{0}\in\mathbb{Z}$ such that for any $a\leq a_{0}$,
$\max_{1\leq k\leq N_{0}}\|\vec{f}_{k}-E_{d,a}\vec{f}_{k}\|_{M_{p}^{t,r}(W)}<\varepsilon.$
Now if $a\leq a_{0}$, then
$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}$
$\leq\|E_{d,a}\vec{f}-E_{d,a}\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|E_{d,a}\vec{f}_{k}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}-\vec{f}\|_{M_{p}^{t,r}(W)}$
$\leq c\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\varepsilon+\varepsilon$
$\lesssim\varepsilon.$
Remark 4.7
Theorem 4.6 is not true for $1<p<t<r=\infty$. Indeed, let $d=1,W\equiv 1$, $0<p<t<r=\infty$. Let $f(x)=|x|^{-n/t}$. Then
$\|f\|_{M^{t,\infty}_{p}}$ $=\sup_{Q\in{\mathcal{D}}}|Q|^{1/t-1/p}\left(\int_{Q}|y|^{-np/t}{\mathrm{d}}y\right)^{1/p}\approx 1.$
Let $f_{j}=\chi_{B(0,2^{j})}f$ for $j\in\mathbb{N}$. Note that $f$ and $f_{j}$ are radial and symmetric functions. Let $x=(x_{1},0,\ldots,0)$ where $x_{1}=2^{j}+2\times 2^{j}$. Let $s=x_{1}/10$ Then $B(x,s)\subset B(0,2^{j})^{c}$. Then
$\|f-f_{j}\|_{M^{t,\infty}_{p}}$
$\geq cs^{-n(1/t-1/p)}\left(\int_{B(x,s)}|y|^{-np/t}{\mathrm{d}}y\right)^{1/p}$
$\geq cs^{-n(1/t-1/p)}\left(\int_{|x|-|s|}^{|x|+|s|}\eta^{-np/t}\eta^{n-1}{\mathrm{d}}\eta\right)^{1/p}$
$=c\left(\frac{|x|}{10}\right)^{-n(1/t-1/p)}\left(\left(1+\frac{1}{10}\right)^{n-np/t}-\left(1-\frac{1}{10}\right)^{n-np/t}\right)^{1/p}|x|^{n(1/t-1/p)}$
$=c>0.$
This shows that the totally bounded set =$\mathcal{F}=\{f\}$ of $M^{t,\infty}_{p}$ is not uniformly vanishes at infinity.
  • Acknowledgement(s).  The corresponding author Jingshi Xu is supported by the National Natural Science Foundation of China (Grant No. 12161022) and the Science and Technology Project of Guangxi (Guike AD23023002).

### v2

@@
-Precompactness in matrix weighted Bourgain-Morrey spaces
-Abstract.
+
+Abstract
 In this paper, we introduce matrix weighted Bourgain-Morrey spaces and obtain two sufficient conditions for precompact sets in matrix weighted Bourgain-Morrey spaces.
@@
 With this result, we obtain the necessity for precompact sets in some matrix weighted Bourgain-Morrey spaces.
+The results are new even for the unweighted Bourgain-Morrey spaces.
+Keywords Kolmogorov-Riesz theorem, matrix weight, precompactness, Bourgain-Morrey space.
+Mathematics Subject Classification Primary 46B50, 46E40, 42B35, 46E30.
 1 Introduction
@@
 The theory of matrix weighted function spaces goes back to [35].
-Indeed, in 1958, Wiener and Masani [32, Section 4] studied the matrix weighted $L^{2}(W)$ for the prediction theory for multivariate stochastic processes.
+Indeed, in 1958, Wiener and Masani [35, Section 4] studied the matrix weighted $L^{2}(W)$ for the prediction theory for multivariate stochastic processes.
 In [32], Treil and Volberg introduced matrix class $\mathcal{A}_{2}$.
@@
 In [7], Bu, Hytönen, Yang, Yuan obtained several new characterizations of $\mathcal{A}_{p,\infty}$-matrix weights.
+In [8], Bu et al.
+studied the inhomogeneous Besov-type and Triebel–Lizorkin-type spaces with the result in [7].
+In [2], Bu et al.
+introduced the matrix weighted Hardy spaces and obtained characterizations of these spaces via maximal function, atom.
+As applications, they established the finite atomic characterization of matrix weighted Hardy spaces and obtained a criterion on the boundedness of sublinear operators from matrix weighted Hardy spaces to any $\gamma$-quasi-Banach space ($\gamma\in(0,1]$).
+The boundedness of Calderón-Zygmund operators on matrix weighted Hardy spaces was also obtained.
+In [23], Li, Yang and Yuan introduced the matrix-weighted Besov-Triebel-Lizorkin spaces with logarithmic smoothness and characterize these spaces via Peetre-type maximal functions.
+In [37, 38], Zhao et al.
+introduced (generalized grand) Besov-Bourgain-Morrey spaces and explored various real-variable properties of these spaces, which are a bridge connecting Bourgain-Morrey spaces with amalgam-type spaces.
+Moreover, some real-variable properties and boundedness of classical operators were studied in their article.
 For many other results on the matrix class $\mathcal{A}_{p}$ and matrix weighted function spaces, we refer the reader to [6, 10, 11, 12, 13, 28, 29, 34].
@@
 Using this result, we obtain the Kolmogorov-Riesz compactness theorem in matrix weighted Bourgain-Morrey spaces.
+These results are new even for the unweighted Bourgain-Morrey spaces.
 Throughout this paper, we let $c,C$ denote constants that are independent of the main parameters involved but whose value may differ from line to line.
-Let =$\mathbb{N}=\{1,2,3,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\ldots\}$.
+Let $\mathbb{N}=\{1,2,3,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\ldots\}$.
 Let $\mathbb{Z}$ be the set of integers.
@@
 We denote by $\mathcal{D}$ the the family of all dyadic cubes in $\mathbb{R}^{n}$, while $\mathcal{D}_{j}$ is the set of all dyadic cubes with $\ell(Q)=2^{-j},j\in\mathbb{Z}$.
-Let $x_{Q}$ be the left corner of $Q\in{\mathcal{D}}$.
+Let $x_{Q}$ be the lower left corner of $Q\in{\mathcal{D}}$.
 For $\lambda>0$, let $\lambda Q$ be the cube with the same center of $Q$ and the edge length $\lambda\ell(Q)$.
@@
 That is, for $1\leq p<q\leq\infty$,
-$|x|_{q}\leq|x|_{p}\leq d^{1/p}|x|_{\infty}\leq d^{1/p}|x|_{q}.$ (2.1)
+$|x|_{q}\leq|x|_{p}\leq d^{1/p}|x|_{\infty}\leq d^{1/p}|x|_{q}.$ (1)
 For $A\in M_{d}(\mathbb{C})$, let
 $\left\|A\right\|:=\sup_{\vec{z}\in\mathbb{C}^{d},\left|\vec{z}\right|=1}\left|A\vec{z}\right|.$
-We say that a matrix $A\in M_{d}(\mathbb{C})$ is positive definite if, for any $\vec{x}\in\mathbb{C}^{d}\backslash\{\vec{0}\}$, $\vec{z}^{*}A\vec{z}>0.$ And a matrix $A\in M_{d}(\mathbb{C})$ is called nonnegative definite if, for any $\vec{x}\in\mathbb{C}^{d}$, $\vec{z}^{*}A\vec{z}\geq 0.$
-From [18, Theorem 4.1.4], any nonnegative definite matrix is always Hermitian. Hence any nonnegative definite matrix is self-adjoint.
+[…]
+From [20, Theorem 4.1.4], any nonnegative definite matrix is always Hermitian. Hence any nonnegative definite matrix is self-adjoint.
 Let $A\in M_{d}(\mathbb{C})$ be a positive definite matrix and have eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$.
-From [18, Theorem 2.5.6(c)], there exists a unitary matrix $U\in M_{d}(\mathbb{C})$ such that
-$A=U\operatorname{diag}\,(\lambda_{1},\ldots,\lambda_{d})U^{*}.$ (2.2)
-Moreover, by [18, Theorem 4.1.8], we find $\{\lambda_{i}\}_{i=1}^{d}\subset(0,\infty)^{d}$. The following definition is based on these conclusions.
-Definition 2.1
+From [20, Theorem 2.5.6(c)], there exists a unitary matrix $U\in M_{d}(\mathbb{C})$ such that
+$A=U\operatorname{diag}\,(\lambda_{1},\ldots,\lambda_{d})U^{*}.$ (2)
+Moreover, by [20, Theorem 4.1.8], we find $\{\lambda_{i}\}_{i=1}^{d}\subset(0,\infty)$. The following definition is based on these conclusions.
+Definition 2.1.
 Let $A\in M_{d}(\mathbb{C})$ be a positive definite matrix with positive eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$. For any $\alpha\in\mathbb{R}$, define
-$A^{\alpha}:=U\operatorname{diag}\left(\lambda_{1}^{\alpha},\ldots,\lambda_{m}^{\alpha}\right)U^{*},$ (2.3)
-where $U$ is the same as in (2.2).
-Remark 2.2
-From [17, p. 408], we obtain that $A^{\alpha}$ is independent of the choices of the order of $\{\lambda_{i}\}_{i=1}^{m}$ and $U$, and hence $A^{\alpha}$ is well defined.
+$A^{\alpha}:=U\operatorname{diag}\left(\lambda_{1}^{\alpha},\ldots,\lambda_{m}^{\alpha}\right)U^{*},$ (3)
+where $U$ is the same as in (2).
+Remark 2.2.
+From [19, p. 408], we obtain that $A^{\alpha}$ is independent of the choices of the order of $\{\lambda_{i}\}_{i=1}^{m}$ and $U$, and hence $A^{\alpha}$ is well defined.
 Now, we recall some concepts of matrix wights.
-Definition 2.3
+Definition 2.3.
 A matrix-valued function $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ is called a matrix weight if $W$ satisfies that
@@
 (ii) for almost every $x\in{\mathbb{R}^{n}}$, $W(x)$ is invertible;
-(iii) the entries of $W$ is locally integrable;
-Definition 2.4
+(iii) the entries of $W$ is locally integrable.
+Definition 2.4.
 Let $p\in(0,\infty)$, $W$ be a matrix weight.
@@
 Then the matrix $A_{E}\in M_{d}(\mathbb{C})$ is called a reducing operator of order $p$ for $W$ if $A_{E}$ is positive definite and, for any $\vec{z}\in\mathbb{C}^{d}$,
-$\left|A_{E}\vec{z}\right|\approx\left(\frac{1}{|E|}\int_{E}\left|W^{\frac{1}{p}}(x)\vec{z}\right|^{p}\,{\mathrm{d}}x\right)^{\frac{1}{p}},$ (2.4)
+$\left|A_{E}\vec{z}\right|\approx\left(\frac{1}{|E|}\int_{E}\left|W^{\frac{1}{p}}(x)\vec{z}\right|^{p}\,{\mathrm{d}}x\right)^{\frac{1}{p}},$ (4)
 where the positive equivalence constants depend only on $d$ and $p$.
-Next we recall the concepts of scalar weight class $A_{p}$ (see [12, Definitions 7.1.1, 7.1.3]) and matrix weight class $\mathcal{A}_{p}$ (see [26] for $1<p<\infty$, [9] for $0<p\leq 1$).
-Definition 2.5
+Next we recall the concepts of scalar weight class $A_{p}$ (see [14, Definitions 7.1.1, 7.1.3]) and matrix weight class $\mathcal{A}_{p}$ (see [27] for $p<\infty, [11$] for $0<p\leq 1$).
+Definition 2.5.
 A weight $\omega$ is a nonnegative locally integrable function on $\mathbb{R}^{n}$} such that $0<\omega(x)<\infty$ for almost all $x\in{\mathbb{R}^{n}}$.
@@
 For $0<p\leq 1$, a matrix weight $W\in\mathcal{A}_{p}(\mathbb{\mathbb{R}}^{n})$ if and only if
-[…]
+$\underset{Q}{\sup}\;\underset{y\in Q}{\mathop{\mathrm{ess\;sup}}}\frac{1}{|Q|}\int_{Q}\|W^{1/p}(x)W^{-1/p}(y)\|^{p}\mathrm{d}x<\infty.$
 We write $\mathcal{A}_{p}:=\mathcal{A}_{p}(\mathbb{\mathbb{R}}^{n})$ for brevity.
-Given any matrix weight $W$ and $0<p<\infty$, there exists (see e.g., [11, Proposition 1.2]) for $p>1$ and [9, p.1237] for $0<p\leq 1)$ a sequence $\{A_{Q}\}_{Q\in\mathcal{D}}$ of positive definite $d\times d$ matrices such that
+Given any matrix weight $W$ and $0<p<\infty$, there exists (see e.g., [13, Proposition 1.2] for $p>1$ and [11, p.1237] for $0<p\leq 1)$ a sequence $\{A_{Q}\}_{Q\in\mathcal{D}}$ of positive definite $d\times d$ matrices such that
 $c_{1}|A_{Q}\vec{y}|\leq\Big(\frac{1}{|Q|}\int_{Q}|W^{1/p}(x)\vec{y}|^{p}\mathrm{d}x\Big)^{1/p}\leq c_{2}|A_{Q}\vec{y}|,$
@@
 In this case, we call $\{A_{Q}\}_{Q\in\mathcal{D}}$ a sequence of reducing operators of order $p$ for $W$.
-Definition 2.6
+Definition 2.6.
 A matrix weight $W$ is called a doubling matrix weight of order $p>0$ if the scalar measures $w_{\vec{y}}(x)=|W^{1/p}(x)\vec{y}|^{p}$, for $\vec{y}\in\mathbb{C}^{d}$, are uniformly doubling: there exists $c>0$ such that for all cubes $Q\subset\mathbb{R}^{n}$ and all $\vec{y}\in\mathbb{C}^{d}$,
-[…]
+$\int_{2Q}w_{\vec{y}}(x)dx\leq c\int_{Q}w_{\vec{y}}(x)\mathrm{d}x.$
 If $c=2^{\beta}$ is the smallest constant for which this inequality holds, we say that $\beta$ is the doubling exponent of $W$.
-From [15, Proposition 2.10], we know that $\beta$ is always not less than $n$.
-In [3], Bu, Hytönen, Yang and Yuan introduced the $\mathcal{A}_{p}$-dimension of matrix weights, which will be used in Theorems 4.5 and 4.6.
-Definition 2.7
+From [17, Proposition 2.10], we know that $\beta$ is always not less than $n$.
+In [3], Bu, Hytönen, Yang and Yuan introduced the $\mathcal{A}_{p}$-dimension of matrix weights, which will be used in Theorems 4.6 and 4.7.
+Definition 2.7.
 Let $0<p<\infty$, $\tilde{d}\in\mathbb{R}$.
 A matrix weight $W$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}$, denoted by $W\in\mathbb{D}_{p,\tilde{d}}(\mathbb{R}^{n},\mathbb{C}^{n})$, if there exists a positive constant $C$ such that for any cube $Q\subset\mathbb{R}^{n}$ and any $i\in\mathbb{N}_{0}$,
-[…]
+$\underset{y\in 2^{i}Q}{\operatorname{ess\;\sup}}\frac{1}{|Q|}\int_{Q}\left\|W^{1/p}(x)W^{-1/p}(y)\right\|^{p}{\mathrm{d}}x\leq C2^{i\tilde{d}},\quad\operatorname{for}\;0<p\leq 1,$ (5)
 or,
-[…]
-where $11/p+1/p^{\prime}=$ 1.
+$\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|2^{i}Q|}\int_{2^{i}Q}\left\|W^{1/p}(x)W^{-1/p}(y)\right\|^{p^{\prime}}{\mathrm{d}}y\right)^{p/p^{\prime}}{\mathrm{d}}x\leq C2^{i\tilde{d}},\quad\operatorname{for}\;1<p<\infty,$
+where $1/p+1/p^{\prime}=1$.
 We denote $\mathbb{D}_{p,\tilde{d}}(\mathbb{R}^{n},\mathbb{C}^{n})$ simply by $\mathbb{D}_{p,\tilde{d}}$.
-Lemma 2.8 (Corollary 2.32, [3])
-Let $0<p<\infty$, let $W\in\mathcal{A}_{p}$ have the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$, and let $\{A_{Q}\}_{\operatorname{cube}Q}$ be the reducing operator of order $p$ for $W$.
-(i) If $1<p<\infty$, let $\widetilde{W}:=W^{-1/(p-1)}$ (which belongs to $\mathcal{A}_{p^{\prime}}$) have the $\mathcal{A}_{p^{\prime}}$-dimension $\tilde{\tilde{d}}$.
+The following lemma says that if $W\in\mathcal{A}_{p}$ for $p\in(0,\infty)$, then $W\in\mathbb{D}_{p,\tilde{d}}(\mathbb{R}^{n},\mathbb{C}^{n})$.
+Lemma 2.8 (Proposition 2.27, [3]).
+Let $p\in(0,\infty)$ and $W\in\mathcal{A}_{p}$. Then there exists $\tilde{d}\in[0,n)$ such that $W$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}$.
+Lemma 2.9 (Corollary 2.32, [3]).
+Let $0<p<\infty$, let $W\in\mathcal{A}_{p}$ with the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$, and let $\{A_{Q}\}_{\operatorname{cube}Q}$ be the reducing operator of order $p$ for $W$.
+(i) If $1<p<\infty$, let $\widetilde{W}:=W^{-1/(p-1)}$ (which belongs to $\mathcal{A}_{p^{\prime}}$) with the $\mathcal{A}_{p^{\prime}}$-dimension $\tilde{\tilde{d}}$.
 Then there exists a positive constant $C$ such that, for any cubes $Q$ and $R$ of $\mathbb{R}^{n}$},
@@
 2.2 Matrix weighted Bourgain-Morrey spaces
-Definition 2.9
+Definition 2.10.
 Let $0<p<\infty$, $d\in\mathbb{N}$ and $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight.
@@
 $\|\vec{f}\chi_{K}\|_{L^{p}(W)}=\left(\int_{K}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}<\infty.$
-Let $\varOmega\subset{\mathbb{R}^{n}}$ be an open set. The space $L^{p}(W,\varOmega)$ collects all measurable functions $\vec{f}:\varOmega\to\mathbb{C}^{d}$ such that
-$\|\vec{f}\|_{L^{p}(W,\varOmega)}=\left(\int_{\varOmega}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}<\infty.$
-Definition 2.10
+Let $\Omega\subset{\mathbb{R}^{n}}$ be an open set. The space $L^{p}(W,\Omega)$ collects all measurable functions $\vec{f}:\Omega\to\mathbb{C}^{d}$ such that
+$\|\vec{f}\|_{L^{p}(W,\Omega)}=\left(\int_{\Omega}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}<\infty.$
+Definition 2.11.
 Let ${\mathcal{D}}=\{Q_{j,k}\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}$ be the standard dyadic system.
@@
 where $W(Q_{j,k})=\int_{Q_{j,k}}\|W(y)\|{\mathrm{d}}y$.
-Let $\{A_{Q}\}_{Q\in{\mathcal{D}}}$ be the reducing operator of order $p$ for $W$. Define $M_{p}^{t,r}(\{A_{Q}\})$ as the set of all $\vec{f}\in L_{\rm loc}^{p}(\{W\})$ such that
-$\|\vec{f}\|_{M_{p}^{t,r}(\{A_{Q}\})}$
-:=$:=\left\|\left\{\left(\|A_{Q_{j,k}}\|^{p}|Q_{j,k}|\right)^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(y)|^{p}{\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$
-Remark 2.11
-Let $\{\vec{e}_{i}\}_{i=1}^{d}$ be the standard unit basis for $\mathbb{C}^{d}$.
-Since $W$ is a matrix weight, then for each $x\in{\mathbb{R}^{n}}$, $W(x)$ is a nonnegative matrix and self-adjoint.
-Hence we have
-$\|W(x)\|$ $=\|W^{2/p}(x)\|^{p/2}\approx[\operatorname{tr}(W^{2/p}(x))]^{p/2}$
-$=\left(\sum_{i=1}^{d}|W^{1/p}(x)\vec{e}_{i}|^{2}\right)^{p/2}\approx\sum_{i=1}^{d}|W^{1/p}(x)\vec{e}_{i}|^{p}.$
-From this and (2.4), we obtain
-$W(Q_{j,k})$ $=\int_{Q_{j,k}}\|W(y)\|{\mathrm{d}}y\approx\int_{Q_{j,k}}\sum_{i=1}^{d}|W^{1/p}(x)\vec{e}_{i}|^{p}{\mathrm{d}}y$
-$\approx\sum_{i=1}^{d}|A_{Q_{j,k}}\vec{e}_{i}|^{p}|Q_{j,k}|\approx\|A_{Q_{j,k}}\|^{p}|Q_{j,k}|.$
+Let $\{A_{Q}\}_{Q\in{\mathcal{D}}}$ be the reducing operator of order $p$ for $W$. Define $M_{p}^{t,r}(\{A_{Q}\})$ as the set of all $\vec{f}\in L_{\rm loc}^{p}(W)$ such that
+$\|\vec{f}\|_{M_{p}^{t,r}(\{A_{Q}\})}:=\left\|\left\{\left(\|A_{Q_{j,k}}\|^{p}|Q_{j,k}|\right)^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(y)|^{p}{\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$
+Remark 2.12.
+By [3, Lemma 2.11] with $M=I_{m}$, we conclude that, for any cube $Q\subset{\mathbb{R}^{n}}$,
+$\|A_{Q}\|^{p}\approx\frac{1}{|Q|}\int_{Q}\left\|W^{1/p}(x)\right\|^{p}{\mathrm{d}}x=\frac{1}{|Q|}\int_{Q}\left\|W(x)\right\|{\mathrm{d}}x=\frac{1}{|Q|}W(Q).$
 Thus $M_{p}^{t,r}(W)$ is same with $M_{p}^{t,r}(\{A_{Q}\})$ in meaning of the equivalent quasi-norms.
 If $d=1,W\equiv 1$, $M_{p}^{t,r}(W)$ is the classical Bourgain-Morrey space $M_{p}^{t,r}$ in [18]. We define the scalar weighted Bourgain-Morrey space $M_{p}^{t,r}(\omega)$ by
-$\|f\|_{M_{p}^{t,r}(\omega)}:=\left\|\left\{(\omega(Q_{j,k}))^{1/t-1/p}\left(\int_{Q_{j,k}}|f(y)|^{p}\omega(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$ (2.6)
+$\|f\|_{M_{p}^{t,r}(\omega)}:=\left\|\left\{(\omega(Q_{j,k}))^{1/t-1/p}\left(\int_{Q_{j,k}}|f(y)|^{p}\omega(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$ (6)
 That is $M_{p}^{t,r}(\omega)$ is the case $d=1$ of matrix weighted Bourgain-Morrey space $M_{p}^{t,r}(W)$.
 The following lemma is proved in [25, 31].
-Lemma 2.12
+Lemma 2.13.
 Let $1\leq p<\infty$ and $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight.
-Let $\varOmega\subset{\mathbb{R}^{n}}$ be an open set.
-Then matrix weighted Lebesgue space $L^{p}(W,\varOmega)$ is a Banach space.
-Theorem 2.13
+Let $\Omega\subset{\mathbb{R}^{n}}$ be an open set.
+Then matrix weighted Lebesgue space $L^{p}(W,\Omega)$ is a Banach space.
+Theorem 2.14.
 Let $1\leq p<\infty$ and $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. Then the space $L^{p}_{\operatorname{loc}}(W)$ is complete.
-  • Proof
+Proof.
 Let $\{\vec{f}_{k}\}_{k=1}^{\infty}$ be a Cauchy sequence in $L^{p}_{\operatorname{loc}}(W)$.
-That is, for any compact set $K$, any $\varepsilon>0$, there exists $N>0$ such that if $j,k>N$, $\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\varepsilon.$ Then for any compact set $K$, $\{\vec{f}_{k}\chi_{K}\}_{k=1}^{\infty}$ is a Cauchy sequence in $L^{p}(W)$.
+That is, for any compact set $K$, any $\epsilon>0$, there exists $N>0$ such that if $j,k>N$, $\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\epsilon.$ Then for any compact set $K$, $\{\vec{f}_{k}\chi_{K}\}_{k=1}^{\infty}$ is a Cauchy sequence in $L^{p}(W)$.
 Since $L^{p}(W)$ is complete (Lemma 2.13), there exists $\vec{f}\chi_{K}$ in $L^{p}(W)$ such that $\vec{f}_{k}\chi_{K}\to\vec{f}\chi_{K}$.
@@
 By the dominated convergence theorem, for $j,k>N$, we obtain
-    $\|(\vec{f}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}=\lim_{j\to\infty}\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\varepsilon.$
-    Hence $\vec{f}_{k}\to\vec{f}$ in $L^{p}_{\operatorname{loc}}(W)$ and the space $L^{p}_{\operatorname{loc}}(W)$ is complete. $\Box$
+$\|(\vec{f}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}=\lim_{j\to\infty}\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\epsilon.$
+Hence $\vec{f}_{k}\to\vec{f}$ in $L^{p}_{\operatorname{loc}}(W)$ and the space $L^{p}_{\operatorname{loc}}(W)$ is complete. ∎
 In what follows, the symbol $\hookrightarrow$ always stands for continuous embedding.
-Proposition 1
+Proposition 2.15.
 Let $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight.
@@
 $L^{t}(W)\hookrightarrow M_{p}^{t,r}(W)\hookrightarrow L_{\operatorname{loc}}^{p}(W).$
-  • Proof
+Proof.
 (i) It comes from $\ell^{r_{1}}\hookrightarrow\ell^{r_{2}}$ since $0<r_{1}<r_{2}\leq\infty$.
@@
 $\leq\left(\frac{1}{W(Q)}\int_{Q}\|W^{1/p_{1}-1/p_{2}}(y)\|^{p_{1}}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{1}}{\mathrm{d}}y\right)^{1/p_{1}}$
-    $\leq\left(\frac{1}{W(Q)}\int_{Q}\|W^{1/p_{1}-1/p_{2}}(y)\|^{p_{1}p_{2}/(p_{2}-p_{1})}{\mathrm{d}}y\right)^{1/p_{1}-1/p_{2}}$
-    ×$\quad\times\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{2}}{\mathrm{d}}y\right)^{1/p_{2}}$
+$\leq\left(\frac{1}{W(Q)}\int_{Q}\|W^{1/p_{1}-1/p_{2}}(y)\|^{p_{1}p_{2}/(p_{2}-p_{1})}{\mathrm{d}}y\right)^{1/p_{1}-1/p_{2}}\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{2}}{\mathrm{d}}y\right)^{1/p_{2}}$
 $=\left(\frac{1}{W(Q)}\int_{Q}\|W(y)\|{\mathrm{d}}y\right)^{1/p_{1}-1/p_{2}}\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{2}}{\mathrm{d}}y\right)^{1/p_{2}}$
@@
 $L^{t}(W)=M_{t}^{t,\infty}(W)\hookrightarrow M_{p}^{t,\infty}(W).$
-    For any compact set $K\subset{\mathbb{R}^{n}}$, there exist at most $2^{n}$ dyadic cubes $Q_{j}$ such that $K\subset\cup_{j=1}^{2n}Q_{j}$. Hence
-    $\|\vec{f}\chi_{K}\|_{L^{p}(W)}$ $\leq\sum_{j=1}^{2^{n}}\frac{|Q_{j}|^{1/t-1/p}}{|Q_{j}|^{1/t-1/p}}\|\vec{f}\chi_{Q_{j}}\|_{L^{p}(W)}$
-    $\leq\|\vec{f}\|_{M_{P}^{t,\infty}(W)}\sum_{j=1}^{2^{n}}\frac{1}{|Q_{j}|^{1/t-1/p}}<\infty.$
+For any compact set $K\subset{\mathbb{R}^{n}}$, there exist at most $2^{n}$ dyadic cubes $Q_{j}$ such that $K\subset\bigcup_{j=1}^{2^{n}}Q_{j}$. Hence
+$\|\vec{f}\chi_{K}\|_{L^{p}(W)}\leq\sum_{j=1}^{2^{n}}\frac{|Q_{j}|^{1/t-1/p}}{|Q_{j}|^{1/t-1/p}}\|\vec{f}\chi_{Q_{j}}\|_{L^{p}(W)}\leq\|\vec{f}\|_{M_{P}^{t,\infty}(W)}\sum_{j=1}^{2^{n}}\frac{1}{|Q_{j}|^{1/t-1/p}}<\infty.$
 This shows that $M_{p}^{t,r}(W)\hookrightarrow L_{\operatorname{loc}}^{p}(W).$
-    $\Box$
-Proposition 2
+Proposition 2.16.
 The matrix weighted Bourgain-Morrey space $M_{p}^{t,r}(W)$ is a Banach space when $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$.
-  • Proof
-    The argument are standard, see, for example, [27, Theorem 2.4]. We only show that $M_{p}^{t,r}(W)$ is complete since others are simple.
+Proof.
+The argument are standard, see, for example, [30, Theorem 2.4]. We only show that $M_{p}^{t,r}(W)$ is complete since others are simple.
 Let $\{\vec{f}_{j}\}_{j=1}^{\infty}$ be a Cauchy sequence in $M_{p}^{t,r}(W)$.
-By Proposition 1, we have that $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is also a Cauchy sequence in $L^{p}_{\operatorname{loc}}(W)$.
+By Proposition 2.15, we have that $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is also a Cauchy sequence in $L^{p}_{\operatorname{loc}}(W)$.
 Since $L^{p}_{\operatorname{loc}}(W)$ is complete, there exists a vector function $\vec{f}$ in $L^{p}_{\operatorname{loc}}(W)$ such that $\vec{f}_{j}\to\vec{f}$.
@@
 $\underset{j\to\infty}{\lim\sup}\|\vec{f}-\vec{f}_{j}\|_{M_{p}^{t,r}(W)}\leq\underset{j\to\infty}{\lim\sup}\left(\underset{k\to\infty}{\lim\inf}\|\vec{f}_{k}-\vec{f}_{j}\|_{M_{p}^{t,r}(W)}\right)=0.$
-Thus, the sequence $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is convergent to $\vec{f}$ in $M_{p}^{t,r}(W)$. $\Box$
+Thus, the sequence $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is convergent to $\vec{f}$ in $M_{p}^{t,r}(W)$. ∎
 Recall that a matrix weight $W$ is almost everywhere invertible. Hereafter, we define $w(x):=\|W^{-1}(x)\|^{-1}$ when a matrix weight $W$ is invertible at $x\in{\mathbb{R}^{n}}$.
-Lemma 2.14 (Proposition 3.2,[8])
+Lemma 2.17 (Proposition 3.2,[10]).
+Let $1\leq p<\infty$.
 For a matrix weight $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$, we have $0<w(x)\leq\|W(x)\|<\infty$ for a.e.
@@
 Furthermore, $W$ satisfies a two weight, degenerate ellipticity condition: for $\xi\in\mathbb{R}^{d}$,
-$w(x)|\xi|^{p}\leq|W^{1/p}(x)\xi|^{p}\leq\|W(x)\||\xi|^{p}.$ (2.7)
-Remark 2.15
-In [8, Proposition 3.2], it is assumed that $\xi\in\mathbb{R}^{d}$ but it also works for $\xi\in\mathbb{C}^{d}$.
-Proposition 3
+$w(x)|\xi|^{p}\leq|W^{1/p}(x)\xi|^{p}\leq\|W(x)\||\xi|^{p}.$ (7)
+Remark 2.18.
+In [10, Proposition 3.2], it is assumed that $\xi\in\mathbb{R}^{d}$ but it also works for $\xi\in\mathbb{C}^{d}$.
+Proposition 2.19.
 Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$.
@@
 Then $\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}=0$ if and only if $\vec{f}(x)=\vec{g}(x)$ a.e..
-  • Proof
-    Clearly, if $\vec{f}(x)=\vec{g}(x)$ a.e., then $\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}=0$. Then we apply Lemma 2.14 to prove the converse. By the degenerate ellipticity condition (2.7), we have
-    00 0 $=\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}$
-    $\geq\left\|\left\{W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|\vec{f}(y)-\vec{g}(y)|^{p}w(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}.$
-    Since $w(y)>0$ a.e., it follows that $\vec{f}(y)-\vec{g}(y)=0$ a.e.. $\Box$
+Proof.
+Clearly, if $\vec{f}(x)=\vec{g}(x)$ a.e., then $\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}=0$. Then we apply Lemma 2.17 to prove the converse. By the degenerate ellipticity condition (7), we have
+$0=\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}\geq\left\|\left\{W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|\vec{f}(y)-\vec{g}(y)|^{p}w(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}.$
+Since $w(y)>0$ a.e., it follows that $\vec{f}(y)-\vec{g}(y)=0$ a.e.. ∎
 3 Sufficient conditions for precompact sets
-Definition 3.1
+Definition 3.1.
 (i) Suppose that $(X,\rho)$ is a metric space.
-Let $A$ be a set of $X$ and let $\varepsilon>0$.
-A set $E\subset X$ is called an $\varepsilon$-net for $A$ if every point $a\in A$, there exists a point $e\in E$ such that $\rho(a,e)<\varepsilon$.
-(ii) A set $A$ is called totally bounded if, for each $\varepsilon>0$, it possesses a finite $\varepsilon$-net.
+Let $A$ be a set of $X$ and let $\epsilon>0$.
+A set $E\subset X$ is called an $\epsilon$-net for $A$ if every point $a\in A$, there exists a point $e\in E$ such that $\rho(a,e)<\epsilon$.
+(ii) A set $A$ is called totally bounded if, for each $\epsilon>0$, it possesses a finite $\epsilon$-net.
 (iii) A subset in a topological space is precompact if its closure is compact.
 It is well known that in a complete metric space, a set is precompact if and only if it is totally bounded.
-Theorem 3.2
+Theorem 3.2.
 Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$.
@@
 (ii) $\mathcal{F}$ is equicontinuous, that is,
-$\lim_{b\to 0}\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in B(0,b)}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}=0.$ (3.1)
+$\lim_{b\to 0}\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in B(0,b)}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}=0.$ (8)
 Here and what follows, $\tau_{y}$ denotes the translation operator: $\tau_{y}\vec{f}(x):=\vec{f}(x-y)$.
-Remark 3.3
+Remark 3.3.
 In this Remark, we will prove that the conditions (i) and (ii) in Theorem 3.2 together imply that the set $\mathcal{F}$ is bounded.
@@
 Now choose $N\geq 1$ such that $Nb=N|h|>2R$. Then $B(Nh,R)\subset B^{c}(0,R)$. Hence
-$\|\vec{f}\|_{M^{t,r}_{p}(W)}$ $\leq\|\vec{f}\chi_{B(0,R)}\|_{M^{t,r}_{p}(W)}+\|\vec{f}\chi_{B^{c}(0,R)}\|_{M^{t,r}_{p}(W)}$
-$\leq N+\|\vec{f}\chi_{B(Nh,R)}\|_{M^{t,r}_{p}(W)}+\|\vec{f}\chi_{B^{c}(0,R)}\|_{M^{t,r}_{p}(W)}\leq N+2.$
+$\|\vec{f}\|_{M^{t,r}_{p}(W)}\leq\|\vec{f}\chi_{B(0,R)}\|_{M^{t,r}_{p}(W)}+\|\vec{f}\chi_{B^{c}(0,R)}\|_{M^{t,r}_{p}(W)}\leq N+\|\vec{f}\chi_{B(Nh,R)}\|_{M^{t,r}_{p}(W)}+\|\vec{f}\chi_{B^{c}(0,R)}\|_{M^{t,r}_{p}(W)}\leq N+2.$
 This proves that $\mathcal{F}$ is bounded.
-  • Proof of Theorem 3.2.
+Now we begin to show Theorem 3.2.
+Proof.
 Assume that $\mathcal{F}\subset M_{p}^{t,r}(W)$ satisfies (i) and (ii).
-Given $\varepsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\varepsilon$-net of $\mathcal{F}$.
+Given $\epsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\epsilon$-net of $\mathcal{F}$.
 Denote by $R_{i}:=[-2^{i},2^{i})^{n}$ for $i\in\mathbb{Z}$.
 Then from condition (i), there exist a positive integer $m$ large enough such that
-    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (3.2)
+$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (9)
 Moreover, by condition (ii), there exists a negative integer $a$ such that
-    $\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R_{a}}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (3.3)
-    There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\cup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$.
+$\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R_{a}}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (10)
+There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\bigcup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$.
 For any $\vec{f}\in\mathcal{F}$ and $x\in{\mathbb{R}^{n}}$, let
-    $\varPhi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
+$\Phi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
 Then for each fixed $x\in{\mathbb{R}^{n}}$, we have
-    $\left|W(x)\left(\vec{f}(x)-\vec{f}_{Q_{j}}\right)\right|\chi_{Q_{j}}(x)$ $=\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}W(x)\left(\vec{f}(x)-\vec{f}(y)\right){\mathrm{d}}y\right|\chi_{Q_{j}}(x)$ (3.4)
-    $\leq\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x).$ (3.5)
-    We split $\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}$ into three parts:
-    $\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
-    $\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
-    +$\quad+\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
-    +$\quad+\left\{\sum_{k<-m}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}_{Q_{j}}\right)\right|\chi_{Q_{j}}(x)$ $=\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right){\mathrm{d}}y\right|\chi_{Q_{j}}(x)$
+$\leq\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x).$ (11)
+We split $\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}$ into three parts:
+$\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$\quad+\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$\quad+\left\{\sum_{k<-m}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
 $=:S_{1}+S_{2}+S_{3}.$
-    We first estimate $S_{1}$. By (3.4), Jensen’s inequality ($p\geq 1$), the Fubini theorem, we have
-    $S_{1}$$\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.6)
-    ×$\quad\times\left.\left(\int_{Q}\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.7)
-    $\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.8)
-    ×$\quad\times\left.\left(\frac{1}{|Q_{j}|}\int_{Q_{j}}\int_{Q}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$ (3.9)
-    =$=\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.10)
-    ×$\quad\times\left.\left(\frac{1}{|Q_{j}|}\int_{Q}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.11)
-    $=2^{-an/p}\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.12)
-    ×$\quad\times\left.\left(\int_{Q}\int_{x-Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.13)
-    where $x-Q_{j}:=\{x-y:y\in Q_{j}\}$. Since $x\in Q\subset Q_{j}$ for some $j\in\{1,2,\ldots,N\}$, we have $x-Q_{j}\subset R_{a}$. Hence by (3.3), we obtain
+We first estimate $S_{1}$. By (3), Jensen’s inequality ($p\geq 1$), the Fubini theorem, we have
+$S_{1}$ $\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\frac{1}{|Q_{j}|}\int_{Q_{j}}\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$
+$=\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\frac{1}{|Q_{j}|}\int_{Q}\int_{Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$=2^{-an/p}\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\int_{Q}\int_{x-Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (12)
+where $x-Q_{j}:=\{x-y:y\in Q_{j}\}$. Since $x\in Q\subset Q_{j}$ for some $j\in\{1,2,\ldots,N\}$, we have $x-Q_{j}\subset R_{a}$. Hence by (10), we obtain
 […]
-    ×$\quad\left.\times\left(\int_{Q}\left|W(x)(\vec{f}(x)-\vec{f}(x-y))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
-    $\leq\varepsilon.$
 As for $S_{2}$, for each $Q\in{\mathcal{D}}_{k}$ where $k=-m,-m+1,\ldots,-a$, there are $2^{(-k-a)n}$ cubes $Q_{j}\in{\mathcal{D}}_{-a}$ such that $\cup Q_{j}=Q$.
 Denote by $Q_{j_{\ell}}$, $\ell=1,2,\ldots,2^{(-k-a)n}$ these cubes.
-Then by (3.4), Jensen’s inequality ($p\geq 1$), the Fubini theorem, we obtain
-    $S_{2}$ =$=\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.14)
-    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j_{\ell}}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.15)
-    $\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.16)
-    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.17)
-    $\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.18)
-    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$ (3.19)
-    $=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.20)
-    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.21)
-    $=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.22)
-    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{x-Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.23)
-    where $x-Q_{j_{\ell}}:=\{x-y:y\in Q_{j_{\ell}}\}$. Note that $x-Q_{j_{\ell}}\subset R_{a}$ when $x\in Q_{j_{\ell}}$. By (3.3), we have
+Then by (3), Jensen’s inequality ($p\geq 1$), and the Fubini theorem, we obtain
+$S_{2}$ $=\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j_{\ell}}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$
+$=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{x-Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (13)
+where $x-Q_{j_{\ell}}:=\{x-y:y\in Q_{j_{\ell}}\}$. Note that $x-Q_{j_{\ell}}\subset R_{a}$ when $x\in Q_{j_{\ell}}$. By (10), we have
 […]
-    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
-    $\leq\varepsilon.$
 As for $S_{3}$, for each set ${\mathcal{D}}_{k}$, there are only $2^{n}$ cubes $Q\in{\mathcal{D}}_{k}$ such that $Q\cap R_{m}\neq\emptyset$.
@@
 Then
-    $S_{3}$ =$=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\int_{Q_{\ell}}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
-    =$=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\sum_{j=1}^{2^{(m-a)n}}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}.$
-    Similarly as $S_{2}$, from (3.4), Jensen’s inequality ($p\geq 1$), the Fubini theorem, and (3.3), we get that $S_{3}\leq\varepsilon$.
-Together with the estimates of $S_{1}$, $S_{2}$, $S_{3}$, we obtain
-    $\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}\leq 3\varepsilon.$ (3.24)
+$S_{3}$ $=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\int_{Q_{\ell}}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
+$=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\sum_{j=1}^{2^{(m-a)n}}\int_{Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}.$
+Similarly as $S_{2}$, from (3), Jensen’s inequality ($p\geq 1$), the Fubini theorem, and (10), we get that $S_{3}\leq\epsilon$. Together with the estimates of $S_{1}$, $S_{2}$, $S_{3}$, we obtain
+$\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}\leq 3\epsilon.$ (14)
 Note that
-    $\|\vec{f}-\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}$ $\leq\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$
-    $=\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}.$
-    Thus via (3.2) and (3.24), we have
-    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq 4\varepsilon.$ (3.25)
-    From (3.25), it suffices to show that $\varPhi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$.
+$\|\vec{f}-\Phi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}=\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}.$
+Thus via (9) and (14), we have
+$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\Phi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq 4\epsilon.$ (15)
+From (15), it suffices to show that $\Phi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$.
 From Remark 3.3, we have
-    $\sup_{\vec{f}\in\mathcal{F}}\|\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq\sup_{\vec{f}\in\mathcal{F}}\|\varPhi(\vec{f})-\vec{f}\|_{M_{p}^{t,r}(W)}+\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
+$\sup_{\vec{f}\in\mathcal{F}}\|\Phi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq\sup_{\vec{f}\in\mathcal{F}}\|\Phi(\vec{f})-\vec{f}\|_{M_{p}^{t,r}(W)}+\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
 Thus, for any $\vec{f}\in\mathcal{F}$,
-    $|W^{1/p}(x)\varPhi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (3.26)
+$|W^{1/p}(x)\Phi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (16)
 Since $W$ is a matrix weight, by (i), (ii) of Definition 2.3, we have for almost everywhere $x\in{\mathbb{R}^{n}}$, $W(x)$ is a positive definite matrix. Therefore, we obtain
-    $|\varPhi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (3.27)
+$|\Phi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (17)
 which implies $|\vec{f}_{Q_{j}}|<\infty$, $j=1,2,\ldots,N$.
-From this and the entries of $W$ is locally integrable, we see that $\varPhi$ is a map from $\mathcal{F}$ to $\mathcal{B}$, a finite dimensional Banach subspace of $M_{p}^{t,r}(W)$.
-Note that $\varPhi(\mathcal{F})\subset\mathcal{B}$ is bounded, and hence is totally bounded.
+From this and the entries of $W$ is locally integrable, we see that $\Phi$ is a map from $\mathcal{F}$ to $\mathcal{B}$, a finite dimensional Banach subspace of $M_{p}^{t,r}(W)$.
+Note that $\Phi(\mathcal{F})\subset\mathcal{B}$ is bounded, and hence is totally bounded.
 The proof of Theorem 3.2 is complete.
-$\Box$
+∎
 Then we give an application in degenerate Bourgain-Morrey-Sobolev spaces with matrix weights.
-Definition 3.4
+Definition 3.4.
 Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$.
@@
 $\|f\|_{M^{t,r}_{p}(\omega)}:=\left\|\left\{\omega(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|f(y)|^{p}\omega(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}},$
-is the norm of the scalar weighted Bourgain-Morrey space in (2.6).
-Lemma 3.5 (p. 262, [33])
+is the norm of the scalar weighted Bourgain-Morrey space in (6).
+Lemma 3.5 (p. 262, [36]).
 A set $\mathcal{F}$ in metric space $X$ is totally bounded if and only if it is Cauchy-precompact, that is, every sequence has a Cauchy subsequence.
-Corollary 3.6
+Corollary 3.6.
 Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$.
@@
 $\lim_{a\to 0}\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in B(0,a)}\|\vec{f}-\tau_{y}\vec{f}\|_{\mathcal{W}^{1,p,t,r}(W)}=0.$
-  • Proof
-    Note that $\mathcal{F}\subset\mathcal{W}^{1,p,t,r}(W)$ satisfies conditions (i)-(ii) if and only if $\mathcal{F}\subset M^{t,r}_{p}(\omega)$ satisfies conditions (i)-(ii) of Theorem 3.2 ($d=1$) and :=$\nabla\mathcal{F}:=\{\nabla f:f\in\mathcal{F}\}\subset M^{t,r}_{p}(W)$ satisfies (i)-(ii) of Theorem 3.2 ($d=n$).
+Proof.
+Note that $\mathcal{F}\subset\mathcal{W}^{1,p,t,r}(W)$ satisfies conditions (i)-(ii) if and only if $\mathcal{F}\subset M^{t,r}_{p}(\omega)$ satisfies conditions (i)-(ii) of Theorem 3.2 ($d=1$) and $\nabla\mathcal{F}:=\{\nabla f:f\in\mathcal{F}\}\subset M^{t,r}_{p}(W)$ satisfies (i)-(ii) of Theorem 3.2 ($d=n$).
 Hence by Theorem 3.2, we obtain that both $\mathcal{F}\subset M^{t,r}_{p}(\omega)$ and $\nabla\mathcal{F}\subset M^{t,r}_{p}(W)$ are totally bounded.
 Then Corollary 3.6 follows from Lemma 3.5.
-$\Box$
+∎
 4 Dyadic average operator conditions for precompact sets
@@
 $\mathcal{F}$ is equicontinuous by means of the dyadic average operator if
-$\lim_{a\to-\infty,a\in\mathbb{Z}}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=0,$ (4.1)
+$\lim_{a\to-\infty,a\in\mathbb{Z}}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=0,$ (18)
 where
-$E_{d,a}\vec{f}(x):=\sum_{Q\in{\mathcal{D}}_{-a}}\frac{\chi_{Q}(x)}{|Q|}\int_{Q}\vec{f}(y){\mathrm{d}}y.$ (4.2)
-Recall that $x_{Q}$ is the left corner of $Q\in{\mathcal{D}}$. Then for $x\in Q\in\mathcal{D}_{-a}$
+$E_{d,a}\vec{f}(x):=\sum_{Q\in{\mathcal{D}}_{-a}}\frac{\chi_{Q}(x)}{|Q|}\int_{Q}\vec{f}(y){\mathrm{d}}y.$ (19)
+Recall that $x_{Q}$ is the lower left corner of $Q\in{\mathcal{D}}$. Then for $x\in Q\in\mathcal{D}_{-a}$
 $E_{d,a}\vec{f}(x)=\frac{\chi_{Q_{-a,0}}(x-x_{Q})}{|Q_{-a,0}|}\int_{Q_{-a,0}}\vec{f}(y-x_{Q}){\mathrm{d}}y.$
-Then we will prove that (3.1) is stronger than (4.1).
-Indeed, suppose that condition (3.1) holds.
-For any $\varepsilon>0$, there exists a cube $R$ with center 00 and side length $02^{a+1}>$ 0 such that
-$\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (4.3)
-If $x,y\in Q\in{\mathcal{D}}_{-a}$, then $x-y\in R.$ By Jensen’s inequality ($p\geq 1$), Hölder’s inequality ($r/p\geq 1$), (4.3), we obtain
-$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|\frac{1}{|Q_{-a,0}|}\right.\right.\right.$
-×$\quad\times\left.\left.\left.\int_{Q_{-a,0}}W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
-$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\right.\right.$
-×$\quad\times\left.\left.\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right)^{1/r}$
-$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\right.$
-×$\quad\times\left.\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}{\mathrm{d}}y\right)^{1/r}$
-$\leq\frac{1}{|Q_{-a,0}|^{1/r}}|Q_{-a,0}|^{1/r}\sup_{y\in R}\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\right.$
-×$\quad\times\left.\left(\int_{Q}\left|W^{1/p}(x)(\vec{f}(x)-\vec{f}(x-y))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
-$\leq\varepsilon,$
-modified when $r=\infty$. Hence we prove that (3.1) is stronger than (4.1).
+Then we will prove that (8) is stronger than (18). Indeed, suppose that condition (8) holds. For any $\epsilon>0$, there exists a cube $R$ with center 0 and side length $2^{a+1}>0$ such that
+$\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (20)
+If $x,y\in Q\in{\mathcal{D}}_{-a}$, then $x-y\in R.$ By Jensen’s inequality ($p\geq 1$), Hölder’s inequality ($r/p\geq 1$), and (20), we obtain
+$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}$
+$=\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
+$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right)^{1/r}$
+$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}{\mathrm{d}}y\right)^{1/r}$
+$\leq\frac{1}{|Q_{-a,0}|^{1/r}}|Q_{-a,0}|^{1/r}\sup_{y\in R}\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)(\vec{f}(x)-\vec{f}(x-y))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}\leq\epsilon,$
+modified when $r=\infty$. Hence we prove that (8) is stronger than (18).
 Replacing the translation operator by the average operator, we have the following result.
-Theorem 4.1
+Theorem 4.1.
 Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$.
@@
 $\lim_{a\to-\infty,a\in\mathbb{Z}}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=0.$
-where $E_{d,a}$ is same as (4.2).
-  • Proof
+where $E_{d,a}$ is same as (19).
+Proof.
 Assume that $\mathcal{F}\subset M_{p}^{t,r}(W)$ satisfies (i)-(iii).
-Given $\varepsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\varepsilon$-net of $\mathcal{F}$.
+Given $\epsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\epsilon$-net of $\mathcal{F}$.
 Denote by $R_{i}:=[-2^{i},2^{i})^{n}$ for $i\in\mathbb{Z}$.
 Then from condition (ii), there exists a positive integer $m$ large enough such that
-    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (4.4)
+$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (21)
 Moreover, by condition (iii), there exists an integer $a<0$ such that
-    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (4.5)
-    There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\cup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$.
+$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (22)
+There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\bigcup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$.
 For any $\vec{f}\in\mathcal{F}$ and $x\in{\mathbb{R}^{n}}$, let
-    $\varPhi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
+$\Phi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
 Note that for $x\in R_{m}$,
-    $\varPhi(\vec{f})(x)=E_{d,a}\vec{f}(x).$
-    Hence via (4.4) and (4.5), we have
-    $\|\vec{f}-\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}$ $\leq\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$ (4.6)
-    $=\|(\vec{f}-E_{d,a}\vec{f})\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$ (4.7)
-    $\leq\|(\vec{f}-E_{d,a}\vec{f})\|_{M_{p}^{t,r}(W)}+\varepsilon$ (4.8)
-    $\leq 2\varepsilon.$ (4.9)
-    From (4.9), it suffices to show that $\varPhi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$.
-And we have proved that $\varPhi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$ in Theorem 3.2.
+$\Phi(\vec{f})(x)=E_{d,a}\vec{f}(x).$
+Hence via (21) and (22), we have
+$\|\vec{f}-\Phi(\vec{f})\|_{M_{p}^{t,r}(W)}$ $\leq\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$
+$=\|(\vec{f}-E_{d,a}\vec{f})\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$
+$\leq\|(\vec{f}-E_{d,a}\vec{f})\|_{M_{p}^{t,r}(W)}+\epsilon\leq 2\epsilon.$ (23)
+From (23), it suffices to show that $\Phi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$.
+And we have proved that $\Phi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$ in Theorem 3.2.
 Thus the proof of Theorem 4.1 is complete.
-$\Box$
-Lemma 4.2 (Lemma 4.5, [8])
+∎
+Lemma 4.2 (Lemma 4.5, [10]).
 Let $1\leq p<\infty$ and $W\in\mathcal{A}_{p}$. Then $\|W\|$ and $\|W^{-1}\|^{-1}$ are scalar $A_{p}$ weights.
@@
 $E_{Q}(\vec{f})(x)=\frac{1}{|Q|}\int_{Q}\vec{f}(y){\mathrm{d}}y.$
-Theorem 4.3
+Theorem 4.3.
 Let $1\leq p<t<r<\infty$ or $1\leq p\leq t<r=\infty$. If $W\in\mathcal{A}_{p}$, then for any $\vec{f}\in M_{p}^{t,r}(W)$,
 $\lim_{\ell(Q)\to 0}\left|E_{Q}\vec{f}(x)-\vec{f}(x)\right|=0\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$
-  • Proof
-    First, for any $\vec{f}=(f_{1},f_{2},\ldots,f_{d})^{T}\in M_{p}^{t,r}(W)$, it suffices to show that $f_{i}\in L^{1}_{\operatorname{loc}}({\mathbb{R}^{n}})$ for each $i=1,2,\ldots,d.$ Since $W\in\mathcal{A}_{p}$, then by (2.3)
+Proof.
+First, for any $\vec{f}=(f_{1},f_{2},\ldots,f_{d})^{T}\in M_{p}^{t,r}(W)$, it suffices to show that $f_{i}\in L^{1}_{\operatorname{loc}}({\mathbb{R}^{n}})$ for each $i=1,2,\ldots,d.$ Since $W\in\mathcal{A}_{p}$, then by (3)
 $|\vec{f}|^{p}=|W^{-1/p}W^{1/p}\vec{f}|^{p}\leq\|W^{-1/p}\|^{p}|W^{1/p}\vec{f}|^{p}=\|W^{-1}\||W^{1/p}\vec{f}|^{p}.$
@@
 By Hölder’s inequality, we have
-    $\int_{Q}|\vec{f}|{\mathrm{d}}x$ $=\int_{Q}|\vec{f}|\|W^{-1}\|^{-1/p}\|W^{-1}\|^{1/p}{\mathrm{d}}x$
-    $\leq\left(\int_{Q}|\vec{f}|^{p}\|W^{-1}\|^{-1}{\mathrm{d}}x\right)^{1/p}\left(\int_{Q}\|W^{-1}\|^{p^{\prime}/p}{\mathrm{d}}x\right)^{1/p^{\prime}}$
-    $\leq C_{Q,W,p}\left(\int_{Q}|W^{1/p}\vec{f}|^{p}{\mathrm{d}}x\right)^{1/p}$
-    $\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
+$\int_{Q}|\vec{f}(x)|{\mathrm{d}}x$ $=\int_{Q}|\vec{f}(x)|\|W^{-1}(x)\|^{-1/p}\|W^{-1}(x)\|^{1/p}{\mathrm{d}}x$
+$\leq\left(\int_{Q}|\vec{f}(x)|^{p}\|W^{-1}(x)\|^{-1}{\mathrm{d}}x\right)^{1/p}\left(\int_{Q}\|W^{-1}(x)\|^{p^{\prime}/p}{\mathrm{d}}x\right)^{1/p^{\prime}}$
+$\leq C_{Q,W,p}\left(\int_{Q}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
 Hence, $f_{i}\in L^{1}_{\operatorname{loc}}({\mathbb{R}^{n}})$ for each $i=1,2,\ldots,d$. By the classical Lebesgue differentiation theorem, for each $1\leq i\leq d$, we have
@@
 $\left|E_{Q}\vec{f}(x)-\vec{f}(x)\right|\leq d^{1/2}\max_{i}\left|E_{Q}f_{i}(x)-f_{i}(x)\right|.$
+Thus the proof is finished. ∎
 Next we need the boundedness of dyadic average operator. Given any collection $\mathcal{Q}$ of pairwise disjoint cubes $Q\subset{\mathbb{R}^{n}}$, define the averaging operator $A_{\mathcal{Q}}$ by
 $A_{\mathcal{Q}}\vec{f}(x)=\sum_{Q\in\mathcal{Q}}\frac{\chi_{Q}(x)}{|Q|}\int_{Q}\vec{f}(y){\mathrm{d}}y.$
-Lemma 4.4 (Proposition 4.7,[8])
+Lemma 4.4 (Proposition 4.7,[10]).
 Let $1\leq p<\infty$. Let $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. Then $W\in\mathcal{A}_{p}$ if and only it satisfies
 $\|A_{\mathcal{Q}}\vec{f}\|_{L^{p}(W)}\leq C_{n,d,p,W}\|\vec{f}\|_{L^{p}(W)}.$
-Theorem 4.5
-Let $W\in\mathcal{A}_{p}$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$ with the doubling exponent $\beta$.
-(i) Let $1\leq p<t<r<\infty$ If $(\tilde{d}-n+\beta)r/p+n-\beta r/t<0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
-(ii) Let $1\leq p\leq t<r=\infty$ If $\tilde{d}/p-n/p-\beta(1/t-1/p)\leq 0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
-  • Proof
+Remark 4.5.
+In the sequel, let $p,p^{\prime},W,\tilde{d},\widetilde{W},\tilde{\tilde{d}}$ have the same meaning as in Lemma 2.9. Let
+$\tilde{\beta}:=\begin{cases}n,&\operatorname{if}\;p=1,\\ n+\tilde{\tilde{d}}p/p^{\prime},&\operatorname{if}\;1<p<\infty.\end{cases}$ (24)
+Then we claim that for $j+a\in\mathbb{N}_{0}$,
+$W((j+a)_{\operatorname{pa}}Q)\lesssim\begin{cases}2^{(j+a)n}W(Q),&\operatorname{if}p=1,\\ 2^{(j+a)(n+\tilde{\tilde{d}}p/p^{\prime})}W(Q),&\operatorname{if}1<p<\infty.\end{cases}$
+Indeed, when $p=1$, by Lemma 2.9, we have
+$W((j+a)_{\operatorname{pa}}Q)\approx|(j+a)_{\operatorname{pa}}Q|\|A_{(j+a)_{\operatorname{pa}}Q}\|\lesssim 2^{(j+a)n}|Q|\|A_{Q}\|\approx 2^{(j+a)n}W(Q).$
+When $p\in(1,\infty)$, by Lemma 2.9, we obtain
+$W((j+a)_{\operatorname{pa}}Q)$ $\approx|(j+a)_{\operatorname{pa}}Q|\|A_{(j+a)_{\operatorname{pa}}Q}\|^{p}\lesssim 2^{(j+a)n}|Q|2^{(j+a)\tilde{\tilde{d}}p/p^{\prime}}\|A_{Q}\|^{p}\approx 2^{(j+a)(n+\tilde{\tilde{d}}p/p^{\prime})}W(Q).$
+Thus we prove the claim.
+Now we show this estimate is better than $W((j+a)_{\operatorname{pa}}Q)\lesssim 2^{(j+a)\beta}W(Q)$.
+Indeed, when $p=1$, it is obvious since $\beta\geq n$.
+If $p>1$, from [3, Lemma 2.11 and Corollary 2.16], for $i\in\mathbb{N}_{0}$, we deduce that
+$\|A_{i_{\operatorname{pa}}Q}A_{Q}^{-1}\|^{p}$ $\approx\frac{1}{|i_{\operatorname{pa}}Q|}\int_{i_{\operatorname{pa}}Q}\left\|W^{1/p}(x)A_{Q}^{-1}\right\|^{p}{\mathrm{d}}x$
+$\approx\frac{1}{|i_{\operatorname{pa}}Q|}\int_{i_{\operatorname{pa}}Q}\left(\frac{1}{|Q|}\int_{Q}\left\|W^{1/p}(x)W^{-1/p}(y)\right\|^{p^{\prime}}{\mathrm{d}}y\right)^{p/p^{\prime}}{\mathrm{d}}x.$
+From [12, Lemma 2.2], we have $\|A_{i_{\operatorname{pa}}Q}A_{Q}^{-1}\|^{p}\lesssim 2^{i(\beta-n)}$. Hence
+$\frac{1}{|i_{\operatorname{pa}}Q|}\int_{i_{\operatorname{pa}}Q}\left(\frac{1}{|Q|}\int_{Q}\left\|W^{1/p}(x)W^{-1/p}(y)\right\|^{p^{\prime}}{\mathrm{d}}y\right)^{p/p^{\prime}}{\mathrm{d}}x\lesssim 2^{i(\beta-n)}.$
+This, together with [3, Proposition 2.28(ii)], further implies that $W^{-1/(p-1)}\in\mathcal{A}_{p^{\prime}}$ has the $\mathcal{A}_{p^{\prime}}$-dimension $\tilde{\tilde{d}}:=(\beta-n)/(p-1)$ and hence
+$n+\tilde{\tilde{d}}p/p^{\prime}=\beta.$
+Hence we show this estimate is better than $W((j+a)_{\operatorname{pa}}Q)\lesssim 2^{(j+a)\beta}W(Q)$.
+Theorem 4.6.
+Let $W\in\mathcal{A}_{p}$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$. Let $\tilde{\beta}$ be the same with (24).
+(i) Let $1\leq p<t<r<\infty$ .
+If $nr/p+\tilde{d}r/p+n-\tilde{\beta}r(1/t-1/p)<0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
+(ii) Let $1\leq p\leq t<r=\infty$ .
+If $\tilde{d}/p-n/p-\tilde{\beta}(1/t-1/p)\leq 0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
+Proof.
 (i) For a fixed $a\in\mathbb{Z}$, we have
 $\|E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}$ $\leq\left(\sum_{j\leq-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
-    $+\left(\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
-    $=:S_{1}+S_{2}.$
+$+\left(\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}=:S_{1}+S_{2}.$
 By Lemma 4.4, we obtain
-    $S_{1}$ $\leq C_{n,d,p,W}\left(\sum_{j\leq-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
-    $\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}.$
-    When $j>-a$, we denote by $(j+a)_{\operatorname{pa}}Q\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q\in{\mathcal{D}}_{j}$.
+$S_{1}$ $\leq C_{n,d,p,W}\left(\sum_{j\leq-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}.$
+When j>aj>-a, we denote by $(j+a)_{\operatorname{pa}}Q\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q\in{\mathcal{D}}_{j}$.
 Let $\{A_{Q}\}_{Q\in{\mathcal{D}}}$ be the reducing operator of order $p$ for $W$.
 By Lemmas 2.9, we have
-    $\|A_{Q}A_{(j+a)_{\operatorname{pa}}Q}^{-1}\|\leq c2^{(j+a)\tilde{d}/p}.$ (4.10)
-    In $S_{2}$, using (2.5) and (4.10) , we have
+$\|A_{Q}A_{(j+a)_{\operatorname{pa}}Q}^{-1}\|\leq c2^{(j+a)\tilde{d}/p}.$ (25)
+In $S_{2}$, using (5) and (25) , we have
 $S_{2}^{r}$ $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
-    $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}$
-    ×$\quad\times\left(\frac{|Q|}{|Q|}\int_{Q}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}$
+$=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\frac{|Q|}{|Q|}\int_{Q}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}$
 $\approx\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}\left(\left|A_{Q}\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|\right)^{r}$
-    $\lesssim\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}$
-    ×$\quad\times\left(2^{(j+a)\tilde{d}/p}\left|A_{(j+a)_{\operatorname{pa}}Q}\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|\right)^{r}$
-    $\approx\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}2^{(j+a)\tilde{d}r/p}$
-    ×$\quad\times\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y|^{p}{\mathrm{d}}x\right)^{r/p}$
-    $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}$
-    ×$\quad\times\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
-    $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}$
-    ×$\quad\times\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}.$
-    Remark that given $S\in{\mathcal{D}}_{-a}$, there are $2^{(j+a)n}$ cubes $R$ such that $(j+a)R=S$. Hence, by Lemma 4.4 and $(\tilde{d}-n+\beta)r/p+n-\beta r/t<0$, we have
-    $\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}$
-    ×$\times\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
-    $=\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}$
-    ×$\quad\times\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
-    $\leq\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}2^{(j+a)n}2^{-(j+a)\beta(r/t-r/p)}$
-    ×$\quad\times\sum_{S\in\mathcal{D}_{-a}}W(S)^{r/t-r/p}\left(\int_{S}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
-    $\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}^{r},$
-    where in the second step we use the following fact. Since $W((j+a)_{\operatorname{pa}}Q)\leq 2^{(j+a)\beta}W(Q)$ and $01/t-1/p<$ 0, we have
-    $W(Q)^{1/t-1/p}\leq 2^{-(j+a)\beta(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p}.$
+$\lesssim\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}\left(2^{(j+a)\tilde{d}/p}\left|A_{(j+a)_{\operatorname{pa}}Q}\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|\right)^{r}$
+$\approx\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}2^{(j+a)\tilde{d}r/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y|^{p}{\mathrm{d}}x\right)^{r/p}$
+$=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
+$=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}.$
+Remark that given $S\in{\mathcal{D}}_{-a}$, there are $2^{(j+a)n}$ cubes $R$ such that $(j+a)_{\operatorname{pa}}R=S$. From Remark 4.5 and 1/t1/p<01/t-1/p<0, we have
+$W(Q)^{1/t-1/p}\lesssim\begin{cases}2^{-(j+a)n(1/t-1)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1},&\operatorname{if}p=1,\\ 2^{-(j+a)(n+\tilde{\tilde{d}}p/p^{\prime})(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p},&\operatorname{if}1<p<\infty.\end{cases}$
+That is $W(Q)^{1/t-1/p}\lesssim 2^{-(j+a)\tilde{\beta}(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p}$. Hence, by Lemma 4.4 and $nr/p+\tilde{d}r/p+n-\tilde{\beta}r(1/t-1/p)<0$, we have
+$\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
+$=\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
+$\leq\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}2^{(j+a)n}2^{-(j+a)\tilde{\beta}(r/t-r/p)}\sum_{S\in\mathcal{D}_{-a}}W(S)^{r/t-r/p}\left(\int_{S}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
+$\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}^{r}.$
 (ii) Fix $Q_{j,k}\in{\mathcal{D}}$. If $j\leq-a$, then by Lemma 4.4, we have
-$W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}$
-$\lesssim W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}.$
-If $j>-a$, we denote by $(j+a)_{\operatorname{pa}}Q_{j,k}\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q_{j,k}\in{\mathcal{D}}_{j}$.
+$W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}\lesssim W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}.$
+If j>aj>-a, we denote by $(j+a)_{\operatorname{pa}}Q_{j,k}\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q_{j,k}\in{\mathcal{D}}_{j}$.
 $W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}$
@@
 $\lesssim W(Q_{j,k})^{1/t-1/p}|Q_{j,k}|^{1/p}2^{(j+a)\tilde{d}/p}\left|A_{(j+a)_{\operatorname{pa}}Q_{j,k}}\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\vec{f}(y){\mathrm{d}}y\right|$
-$\approx W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\right.$
-×$\quad\times\left.\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{1/p}.$
-Then by Lemma 4.4 and $\tilde{d}/p-n/p-\beta(1/t-1/p)\leq 0$, we obtain
+$\approx W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{1/p}.$
+Then by Lemma 4.4 and $\tilde{d}/p-n/p-\tilde{\beta}(1/t-1/p)\leq 0$, we obtain
 $W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}$
-$\lesssim W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}$
-×$\quad\times\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
+$\lesssim W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
 $=W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}2^{-an/p}\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
-$\lesssim 2^{(j+a)(\tilde{d}/p-n/p)}2^{-(j+a)\beta(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}$
-×$\quad\times\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
-$\lesssim W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
-where in the third step we use the following fact. Since $W((j+a)_{\operatorname{pa}}Q)\leq 2^{(j+a)\beta}W(Q)$ and $1/t-1/p\leq 0$, we have
-$W(Q)^{1/t-1/p}\leq 2^{-(j+a)\beta(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p}.$
+$\lesssim 2^{(j+a)(\tilde{d}/p-n/p)}2^{-(j+a)\tilde{\beta}(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
+$\lesssim W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}.$
 Finally, take the supremum over the cubes $Q_{j,k}\in D$, we obtain
 $\|E_{d,a}\vec{f}\|_{M_{p}^{t,\infty}(W)}\lesssim\|\vec{f}\|_{M_{p}^{t,\infty}(W)}.$
+Hence we finish the proof. ∎
 Finally, we have the following Kolmogorov-Riesz compactness theorem for matrix weighted Bourgain-Morrey spaces.
-Theorem 4.6
-Let $1\leq p<t<r<\infty$ Let $W\in\mathcal{A}_{p}$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$ with the doubling exponent $\beta$.
-Let $(\tilde{d}-n+\beta)r/p+n-\beta r/t<0$.
+Theorem 4.7.
+Let $1\leq p<t<r<\infty$ .
+Let $W\in\mathcal{A}_{p}$ with the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$.
+Let $\tilde{\beta}$ be the same with (24).
+Let $nr/p+\tilde{d}r/p+n-\tilde{\beta}r(1/t-1/p)<0$.
 A subset $\mathcal{F}$ of $M_{p}^{t,r}(W)$ is totally bounded if and only if the following conditions hold:
@@
 $\lim_{a\to-\infty,a\in\mathbb{Z}}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=0.$
-where $E_{d,a}$ is same as (4.2).
-  • Proof
+where $E_{d,a}$ is same as (19).
+Proof.
 The sufficiency is due to Theorem 4.1.
@@
 Assume that $\mathcal{F}$ of $M_{p}^{t,r}(W)$ is totally bounded.
-For any given $\varepsilon>0$, there exists $\{\vec{f}_{k}\}_{k=1}^{N_{0}}\subset\mathcal{F}$ such that $\{\vec{f}_{k}\}_{k=1}^{N_{0}}$ is an $\varepsilon$-net of $\mathcal{F}$, that is, for any $\vec{f}\in\mathcal{F}$, there exists $\vec{f}_{k},k\in\{1,2,\ldots,N_{0}\}$ such that $\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}<\varepsilon$.
+For any given $\epsilon>0$, there exists $\{\vec{f}_{k}\}_{k=1}^{N_{0}}\subset\mathcal{F}$ such that $\{\vec{f}_{k}\}_{k=1}^{N_{0}}$ is an $\epsilon$-net of $\mathcal{F}$, that is, for any $\vec{f}\in\mathcal{F}$, there exists $\vec{f}_{k},k\in\{1,2,\ldots,N_{0}\}$ such that $\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}<\epsilon$.
 Clearly, (i) is true.
-    For any $\varepsilon>0$ and $k\in\{1,2,\ldots,N_{0}\}$, since $r<\infty$, there exists $R_{k}>0$ such that
-    $\|\vec{f}_{k}\chi_{B^{c}(0,R_{k})}\|_{M_{p}^{t,r}(W)}<\varepsilon.$
-    Taking $R=\max_{k\in{1,\ldots,N_{0}}}\{R_{k}\}$, we have $\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ Then for any $\vec{f}\in\mathcal{F}$,
-    $\|\vec{f}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}\leq\|(\vec{f}-\vec{f}_{k})\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<2\varepsilon.$
+For any $\epsilon>0$ and $k\in\{1,2,\ldots,N_{0}\}$, since $r<\infty$, there exists $R_{k}>0$ such that
+$\|\vec{f}_{k}\chi_{B^{c}(0,R_{k})}\|_{M_{p}^{t,r}(W)}<\epsilon.$
+Taking $R=\max_{k\in\{1,\ldots,N_{0}\}}\{R_{k}\}$, we have $\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<\epsilon.$ Then for any $\vec{f}\in\mathcal{F}$,
+$\|\vec{f}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}\leq\|(\vec{f}-\vec{f}_{k})\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<2\epsilon.$
 Thus we prove that
@@
 Thus using the dominated convergence theorem to obtain that there exists $a_{0}\in\mathbb{Z}$ such that for any $a\leq a_{0}$,
-$\max_{1\leq k\leq N_{0}}\|\vec{f}_{k}-E_{d,a}\vec{f}_{k}\|_{M_{p}^{t,r}(W)}<\varepsilon.$
+$\max_{1\leq k\leq N_{0}}\|\vec{f}_{k}-E_{d,a}\vec{f}_{k}\|_{M_{p}^{t,r}(W)}<\epsilon.$
 Now if $a\leq a_{0}$, then
-$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}$
-$\leq\|E_{d,a}\vec{f}-E_{d,a}\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|E_{d,a}\vec{f}_{k}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}-\vec{f}\|_{M_{p}^{t,r}(W)}$
-$\leq c\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\varepsilon+\varepsilon$
-$\lesssim\varepsilon.$
-Remark 4.7
+$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}$ $\leq\|E_{d,a}\vec{f}-E_{d,a}\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|E_{d,a}\vec{f}_{k}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}-\vec{f}\|_{M_{p}^{t,r}(W)}$
+$\leq c\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\epsilon+\epsilon\lesssim\epsilon.$
+Then the proof is complete. ∎
+Remark 4.8.
 Theorem 4.7 is not true for $1<p<t<r=\infty$. Indeed, let $d=1,W\equiv 1$, $0<p<t<r=\infty$. Let $f(x)=|x|^{-n/t}$. Then
@@
 Let $x=(x_{1},0,\ldots,0)$ where $x_{1}=2^{j}+2\times 2^{j}$.
-Let $s=x_{1}/10$ Then $B(x,s)\subset B(0,2^{j})^{c}$.
+Let $s=x_{1}/10$ .
+Then $B(x,s)\subset B(0,2^{j})^{c}$.
 Then
-$\|f-f_{j}\|_{M^{t,\infty}_{p}}$
-$\geq cs^{-n(1/t-1/p)}\left(\int_{B(x,s)}|y|^{-np/t}{\mathrm{d}}y\right)^{1/p}$
-$\geq cs^{-n(1/t-1/p)}\left(\int_{|x|-|s|}^{|x|+|s|}\eta^{-np/t}\eta^{n-1}{\mathrm{d}}\eta\right)^{1/p}$
+$\|f-f_{j}\|_{M^{t,\infty}_{p}}$ $\geq cs^{-n(1/t-1/p)}\left(\int_{B(x,s)}|y|^{-np/t}{\mathrm{d}}y\right)^{1/p}\geq cs^{-n(1/t-1/p)}\left(\int_{|x|-|s|}^{|x|+|s|}\eta^{-np/t}\eta^{n-1}{\mathrm{d}}\eta\right)^{1/p}$
 $=c\left(\frac{|x|}{10}\right)^{-n(1/t-1/p)}\left(\left(1+\frac{1}{10}\right)^{n-np/t}-\left(1-\frac{1}{10}\right)^{n-np/t}\right)^{1/p}|x|^{n(1/t-1/p)}$
 $=c>0.$
-This shows that the totally bounded set =$\mathcal{F}=\{f\}$ of $M^{t,\infty}_{p}$ is not uniformly vanishes at infinity.
-  • Acknowledgement(s).
-The corresponding author Jingshi Xu is supported by the National Natural Science Foundation of China (Grant No.
-12161022) and the Science and Technology Project of Guangxi (Guike AD23023002).
+This shows that the totally bounded set $\mathcal{F}=\{f\}$ of $M^{t,\infty}_{p}$ is not uniformly vanishes at infinity.